### v1

# Evaluating radiation impact on transmon qubits in above and underground facilities

Abstract
Superconducting qubits can be sensitive to abrupt energy deposits caused by cosmic rays and ambient radioactivity. Previous studies have focused on understanding possible correlated effects over time and distance due to cosmic rays. In this study, for the first time, we directly compare the response of a transmon qubit measured initially at the Fermilab SQMS above-ground facilities and then at the deep underground Gran Sasso Laboratory (INFN-LNGS, Italy). We observe same average qubit lifetime T1 of roughly 80 microseconds at above and underground facilities. We then apply a fast decay detection protocol and investigate the time structure, sensitivity and relative rates of triggered events due to radiation versus intrinsic noise, comparing above and underground performance of several high-coherence qubits. Using gamma sources of variable activity we calibrate the response of the qubit to different levels of radiation in an environment with minimal background radiation. Results indicate that qubits respond to a strong gamma source and it is possible to detect particle impacts. However, when comparing above and underground results, we do not observe a difference in radiation induced-like events for these sapphire and niobium-based transmon qubits. We conclude that the majority of these events are not radiation related and to be attributed to other noise sources which by far dominate single qubit errors in modern transmon qubits.
In recent decades, superconducting circuits have emerged as a leading technology explored in the fields of quantum information [1, 2, 3, 4, 5, 6], photon and particle detectors [7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. In the past years, several research groups have started investigating the effect of particle interactions on the performance of quantum bits (qubits) [17, 18]. There are many studies investigating these potential effects in superconducting quantum devices of various geometries and materials. Some early work pointed to the lifetime of a transmon qubit potentially decreasing when exposed to a strong radioactive source [19]. Moreover, the effect of high energy deposits can rapidly propagate from the initial location of deposition, affecting multiple qubits and potentially leading to a general chip failure [20, 21]. C.D. Wilen et al.[22] observed that the rate of charge jumps experienced by an array of qubits (operated as electrometers) was consistent with the rate expected from interactions due to ambient radioactivity and cosmic rays. Thorbeck et al. [23] did not observe significant radiation effects, but concluded that the impacts due to radioactivity may induce frequency jumps in a dominant loss mechanism, namely two-level systems (TLS) [24, 25, 26, 27, 28, 29, 30, 31], possibly affecting qubit stability.
Other studies have shown that suppressing radioactivity can enhance the performance of superconducting circuits. Operating the device in a radio-pure facility has allowed to significantly improve the locking of a fluxonium qubit at the sweet spot [32] and enhanced the internal quality factor of superconducting resonators [33]. The latter publication also enabled the determination that the substrate on which qubits are deposited, rather than the qubits themselves, is the primary target for particle interactions.
These findings have motivated efforts in further directions aimed at studying and mitigating possible effects of radioactivity. Promising results have been achieved by equipping the chip with “traps” designed to absorb phonons produced (also) by radioactive interactions [34, 35, 36] and using gap engineering [18]. Some groups are also focusing on the components used to construct and operate qubits [37, 38], as it has been demonstrated that some of them, such as the Printed Circuit Boards (PCBs), can produce a sizeable rate of events in the qubit chip [39]. Furthermore, it is also worth mentioning recent proposals of equipping qubits with a device specialised in recognizing particle interactions [40] or with external muon detectors [41, 21, 42].
In this work, for the first time, we study charge-insensitive transmon qubit devices in two laboratories with very distinct radiation environments, as a step of the Superconducting Quantum Materials and Systems (SQMS) Center Round Robin project [39], which aims at studying identical qubits measured at multiple test sites around the world. We first compare average relaxation times measured over several hours, and then employ a fast decay detection protocol to identify reductions in relaxation time at rates faster than 10 kHz and analyze time-bin correlations to determine the characteristics of the events. We study the response of the qubit in a controlled environment, by exposing it to gamma-ray sources of increasing activity. Using the same protocol and analysis strategy, we investigate similar qubits hosted above ground in a typical laboratory environment, to compare effect of radiation on transmon qubits performance and behavior, and estimate timescales at which radiation may become concerning for quantum computational errors.
I Experimental locations
To perform the comparative study, we operated similar transmon chips (charge-insensitive) in two different laboratories: the Quantum Garage of the SQMS Center at Fermilab (FNAL, Illinois, US), and the Ieti Facility located in the deep underground Gran Sasso Laboratory of INFN (INFN-LNGS, L’Aquila, Italy). The chip consisted of a 432 m thick, 7.57.5 mm2 wide HEMEX grade Sapphire substrate. Eight Niobium transmons with different geometries were deposited on the chip (see Supplementary section A); their surface was capped using a 10 nm thick layer of Gold (measured at both sites) or Tantalum (measured at FNAL) to reduce losses due to the formation of Nb2O5 [43]. The capping technique allowed to achieve excellent T1 of many tens and up to hundreds of microseconds, which is important to study the change in qubit’s lifetime at microsecond timescales.
We characterized several qubits of these chips operated at FNAL and at INFN-LNGS, and performed comparisons for either an identical qubit or qubits of similar geometries and materials employed. The detailed parameters of these devices are provided in the Supplementary Section A. Additionally, it is noteworthy that the entire substrate is covered in Niobium creating a ground plane; this may act as a phonon trap, diminishing the phonon propagation and absorption in the active part of the qubit [44, 45].
To predict the rate of impacts due to ambient radiation, we conducted a Monte Carlo simulation using the GEANT4 framework [46] developed in Ref. [39]. We anticipate the total rate of radioactive events interacting with the chip at FNAL and INFN-LNGS to be events/sec and events/sec, respectively. Table 1 provides a summary of the simulation results, with detailed information provided in the Supplementary Section C.
Source | FNAL | INFN-LNGS |
[events/sec] | [events/sec] | |
Lab -rays | ||
Muons | ||
Setup | ||
Total |
At FNAL (see Fig. 1(a)), the rate is primarily dominated by rays produced by the decay of naturally radioactive isotopes present in the laboratory environment. However, muons generated by cosmic rays also contribute significantly to the overall rate. In contrast, at INFN-LNGS (see Fig. 1(a)), muons are suppressed by six orders of magnitude thanks to the 1.4 km thick rock overburden. Moreover, the radioactivity of the laboratory is substantially reduced by copper and lead shields positioned both inside and around the refrigerator. The last source of radioactivity arises from the materials constituting the qubit chip and its nearby readout components. In a previous study conducted within the SQMS Round Robin project [39], we measured the radioactive content of each component and determined that cables, connectors, amplifiers, circulators, etc., have a negligible impact on the total event rate. The only exception is the PCB, which is typically highly contaminated and located in close proximity to the chip. In this setup, the PCBs are responsible for the whole counting rate reported in the line “Setup” of Table 1. While this rate is negligible at FNAL, it becomes the dominant source of radioactivity at INFN-LNGS, where all external sources of radioactivity are suppressed.
Furthermore, at INFN-LNGS, we utilize Thorium radioactive sources to increase the event rate up to one event every two seconds and investigate the chip behaviour in a “high”-radiation environment.
II Effect of radiation impact
When a radioactive particle impinges on the qubit chip, it releases energy. Muons create long tracks across the chip, whereas rays interact through photo-electric absorption or (mainly via) Compton scattering, producing short-track electrons. In the absence of an electric field, the thousands of charges created along the ionizing track recombine into phonons, which diffuse throughout the chip [22]. As illustrated in Fig. 2(a), when phonons interact with a superconductor, they can break Cooper pairs into quasiparticles. Subsequently, when one or more quasiparticles tunnel through the Josephson Junction, the qubit looses its energy and quickly decays to its ground state [35] with a timescale of 0.16-1.6 s [47, 48]. Since the metallic pads of the transmon host a large number of quasiparticles that take some time to dissipate through normal metal (PCB leads) or lower-energy phonon scattering, these tunneling events can continue to occur, causing the qubit to suffer from strong relaxation for some time. Previous studies claim this low T1 periods to last from one to several tens of milli-seconds [21, 20]. We investigated these type of events in our studies.
We first performed standard T1 experiments on both underground and above ground locations on the identical qubit Q1 and monitored the relaxation times for close to an hour. These experiments were slow and each data point took about 50 seconds to complete. The results, plotted in Fig. 1(b), show similar mean values and fluctuations, typical for transmon qubits [43], demonstrating that radiation has no significant effect at current state-of-the-art transmon qubit lifetimes of 100 microseconds. No sudden drop in T1 values was observed. This was expected as the standard T1 experiments, due to its averaging nature, are not sensitive to fluctuations at milli-second timescales. In order to study possible events at such a time resolution we employed a protocol, tailored for single qubits, utilizing fast decay measurements [20, 18].
Datasets | acquisition time | live-time |
[min] | [min] | |
Shielded chip | 676.0 | 28.2 |
Shielded chip (2) | – | 124.0 |
44 kBq Th Source | 619.5 | 114.0 |
76 kBq Th Source | 111.7 | 46.3 |
125 kBq Th Source | 60.9 | 28.2 |
161 kBq Th Source | 60.9 | 54.3 |
FNAL Q2 | 234.5 | 104.0 |
FNAL Q3 | 449.7 | 141.9 |
FNAL Q4 | 242.7 | 25.9 |
The fast detection protocol illustrated in Fig. 2(b) begins by resetting the qubit in its excited state . This is accomplished by initially measuring the qubit and applying a conditional -pulse if the qubit is found to be in state via fast feedback. Following a waiting period lasting a few microseconds, the qubit is measured again to determine whether it is in state or . Given a sufficiently long qubit T1, the qubit is highly likely to be found in state . Theoretically, the qubit relaxes to state only a small fraction of the time, given by . However, a radiation event significantly reduces the T1, making it highly probable for the qubit to be detected in state . The sequence is then repeated after another short period, allowing the chip to cool down. Since the effect of quasiparticle generation persists longer than the time taken for one sequence, an event due to radiation typically results in repeated detection in state , as depicted in Fig. 2(c).
In our measurements, we kept the waiting period fixed to 5 s while other parameters like -pulse and readout pulse lengths were similar and reported in the Supplementary Section A. The cooldown period is chosen to ensure that the transmon does not exhibit excessive population in the second excited state . Nevertheless, a simultaneous fit of the , and the second excited state showed that there were many moments in which the (and sometimes even the third excited state ) states were populated. To ensure data consistency, we excluded these time intervals. A thorough investigation of this phenomenon will be mandatory for applications in sensing, as this selection significantly limits the measurement’s live-time (see Table 2).
The readout scheme employed is detailed in the Supplementary Section B. The input signal, filtered through attenuators, low-pass filters, and infrared filters, is sent to the qubit; the output signal is fed into cryogenic triple-junction isolators, and then amplified via a cryogenic HEMT followed by a room temperature amplifier. The qubit readout and control are performed through an RFSoC board using the Quantum Instrumentation Control Kit (QICK) [49]. Each measurement generates a complex number (I/Q values) that are post-processed to distinguish between and based on readout calibration.
We acquired the data for several hours containing several hundred million readouts. During the initial analysis, we simply computed the probability of finding a qubit in for a fixed time window and plot as we moved the window (see Fig. 2(c)). A typical record of as a function of time is shown in Fig. 3(a). The magenta color is the mean of the trace while the fluctuations in the trace look steady. A clear spike around 18.2 seconds corresponds to an event with drastic T1 drop. When zoomed in, these spikes show different pulse shapes as plotted in Fig. 3(b). These pulses generally have a sharp rise but the fall times vary between different instances. As depicted in Fig. 1(b), we found pulses of similar time-profile at both locations irrespective of the specific qubit, pointing to the fact that the drop in relaxation time is not tied to a specific transmon. All pulses observed in our studies last less than ten milli-seconds in agreement with the study in Ref. 21 and thus these pulses require multiple qubit measurements within a milli-second for being detected. For example, the pulses in Fig. 1(c) are visible when a one ms time-window is selected for computing and the signal vanishes when a course time-window of 100 ms is chosen. Clearly, a slow measurement scheme will not be able to resolve individual radiation impact events. In other words, standard T1 measurements, due to its slow nature, will not be affected by radiation impacts (happening at intervals of several seconds) as demonstrated earlier.
A simple detection strategy based on triggering pulses on the data by setting a constant threshold, however, suffers from the problem that a large threshold misses many events, whereas, a lower value picks up a lot of noise arising inherently from the qubit measurement. Furthermore, such a noise is not stable in time leading to unexpected variations when setting a simple constant threshold. As shown in Fig. 3(c), the traces occasionally show noisy behavior which do not follow any adjacent spike. These periods of instability can last from sub-seconds to a minute and will affect regular T1 experiments. Even though it appears to be related to TLS scrambling, we currently do not have a clear understanding of the source of this behavior and it will be a subject of future studies. Consequently, to improve the detection efficiency, we develop a modified strategy discussed below.
III Pulse Reconstruction
For the event detection, we developed an approach based on the probability of observing a certain stream of and . Initially, we performed a simultaneous Gaussian fit to the distribution of I/Q of , , and to correct for time instabilities and removed traces in which the qubit turned out to be in the second excited state . The clouds corresponding to and are then rotated in order to be aligned along one quadrature and projected onto the other. Further information can be found in the Supplementary Section D.
We then performed a Gaussian fit to determine the point at which and were maximally separated, enabling us to convert the measured I/Q values to a sequence of 0 (for ) and 1 (for ). This stream of zeros and ones was further processed to construct pulses for event detection. We started with the assumption that, for each measurement, the probability of the qubit remaining in state is equal to . Since, our readouts took several micro-seconds, we added that to the waiting period to accurately compute . Moreover, we found that the qubit relaxations varied by 25-30 over the data acquisition period (see Fig. 1(b)), consistent with the typical behavior of transmon devices [43]. Therefore, was computed using points (corresponding to approximately one minute worth of data) and used to estimate the probability for the qubit to be in state .
We began by picking the first 50 points of the trace and computing the probability for observing the measured counts of zeros and ones according to the Binomial probability distribution. Subsequently, we moved to the next point of the trace and repeated the calculation. Given that the probability for the qubit to be in state was approximately 85–90%, the probability of observing a large number of zeros in a subsequent series of measurements was very small. Finally, we computed the absolute logarithm () of this probability, simply to avoid working with very small numbers. If the probability of observing this series of events was smaller than 1 (), the software trigger was fired, and we saved a pulse with a length of 450 points, 100 of which were acquired before the pulse to monitor the status of the qubit before the interaction. Typical triggered pulses are shown in Fig. 4.
We then computed several quantities to characterize each pulse, including the value of the maximum, the length (defined as the number of points above threshold), the position of the maximum, as well as the value and root-mean-square (RMS) of the baseline before the pulse. These parameters were used to discard non-physical pulses (see orange trace in Fig. 4), such as those occurring on the tail of a preceding pulse or with a maximum too far from the trigger position. Consistency in data selection was ensured by applying the same criteria across all runs.
IV Results and Discussion
The histogram of Max of the selected pulses is shown in Fig. 5. The single qubit measured in the fully shielded facility at LNGS features a rate of events/sec (solid gray trace). This is obtained by using the measurement time of 28 min, which is selected to filter periods with some unwanted excitation of the state. This shorter time causes larger statistical uncertainty. We repeated the analysis including longer intervals (in which state was sometimes excited) in order to increase the statistics and obtained events/sec (hatched gray trace). This factor two discrepancy can be considered as the systematic uncertainty on the result. The observed background event rate is about 2.5 to 5 times higher than the predicted rate from simulations, which is events/sec. This discrepancy suggests the presence of other sources of decoherence that resemble quasiparticle burst events but cannot be ascribed to radioactive sources.
Figure 5 clearly demonstrates that exposing the chip to a radioactive source results in an excess of events. The rate of these events increases linearly with the intensity of the source itself. Next, we compare the triggered and simulated rates for the various Thorium sources in Fig. 6.
The slope parameter p1 of the linear fit in Fig. 6 shows that setting the threshold at 1 probability allows us to reach a detection efficiency of . This result indicates that qubits are sensitive to laboratory gamma radioactivity, despite the fact that gamma rays deposit a lower amount of energy compared to cosmic rays [22, 33]. This is not a significant concern for the development of quantum processors, as minimal passive shielding around the facility can largely mitigate this radioactive source [39, 50, 51]. Conversely, the absence of a high energy threshold for gamma detection presents interesting opportunities for using qubits as particle detectors [52]. While the primary objective of this work was not to operate a qubit as a particle detector, it is evident that optimizing the threshold or trigger algorithm, along with employing more refined pulse reconstruction algorithms (such as matched filter algorithms), could significantly improve detection efficiency.
However, from the linear fit the intercept p0 results in () events/second. This parameter can be interpreted as the expected rate in the absence of radioactive interactions (if the qubit response to the source is fully linear) and exceeds the one measured in the “background” run. Future work will investigate this discrepancy further.
We deepened the study of the effect of radioactivity by investigating a possible activation of two-level-systems (TLS) when deploying the gamma source. As pointed out in Ref. 23, the ionizing effect of radiation on the chip and the consecutive charge recombination could locally affect the distribution of charges and the electric field that they produce, to which TLSs are sensitive. As a result, the TLSs’ frequencies could shift, getting into resonance or out of resonance with the qubit. This would cause a fluctuation of the qubit T1 lasting longer than the recombination time of the quasiparticles produced by an impact.
We analyzed the data obtained from the measurements done at LNGS when exposing the qubit to the 161 kBq Thorium source. We selected pulses with very high maxima (), so to discard possible noise events. For each pulse we calculated the average values of approximately 10 ms before the trigger and tens of milliseconds after it, when the quasiparticle poisoning was already over. A change in the qubit T1 due to TLS dynamics would result in an average value of after the pulse different from what obtained before the trigger.
We made the histogram of for radioactivity impacts and compared it with what obtained by selecting noise events instead (Fig. 7). We didn’t observe significant differences between the two histograms. To double check this result we exposed our qubit to a Cesium source which activity was an order of magnitude higher than the Th source one. Also in that case, no significant differences between radioactivity events and noise events were observed.
This analysis showed that the fluctuations in the qubit T1 correlated to radioactivity impacts were of the same magnitude of what caused by other sources of noise. Given also the high rate of noise events that we observed, we concluded that ionizing radiation does not play a major role in undermining the T1 stability of the qubit we studied.
Finally, we investigated whether radioactive impacts could excite the qubit from to state. To this aim, we performed the same study described above, but with a reversed measurement protocol: initializing the qubit in the state instead of , and monitoring the qubit exposed to a strong gamma source. Interestingly, exposing the qubit to gamma radiation did not result in an excess of events as demonstrated in Fig. 10. This observation is consistent with what was previously reported in Ref. 20, confirming that quasiparticles generated due to radiation do not deliver energy to transmon qubits.
Finally, we conducted a similar fast decay protocol study using multiple “twin-chips” operated at the above ground facilities at SQMS Fermilab. Here qubits are fully unshielded from cosmic and gamma radiation. The results showed consistency with those obtained at LNGS, as we obtained qualitatively similar distributions in all the pulse parameters defined for our study. The distribution of Max in particular, is reported in Figure 8. The histograms do not reveal any distinct behavior of the chip in the measurements conducted deep underground at LNGS compared to those above-ground at FNAL. The protocol applied to three qubits at FNAL under ambient radiation conditions showed varied responses: similar (Q4), reduced (Q2), and increased (Q3) rates when compared to the response of the maximally shielded qubit at LNGS.
These variations observed in the rates for different qubits measured at FNAL compared to the LNGS background one indicate that modern transmons are dominated by other sources of noise and radiation does not play a major role in the drastic drop in T1 at short time scales. This fact has also been corroborated by another recent study [41] where cosmic rays are not found to be a strong contributor to the most frequent correlated error (four qubits).
Even though we did not experimentally observe a measurable effect of environmental radiation induced errors, we can utilize the simulated rates to give a cap estimate for the relevant timescales at which radiation would start to affect quantum computation.
The probability to have at least one event in a time window T is given by , where is the rate of impacts. Using the simulated rates, for environmental radioactivity we find that in a non-shielded laboratory setting a quantum computation would be affected by radioactivity less than 0.1% (1%) times if the computation is completed within approx (). Note that the computation may include multiple runs of an algorithm and not necessarily require to be a single run lasting for T. At a fully shielded environment like LNGS, due to much lower expected rate, we have instead (1%) if (2.5 s). The available computation period increases if higher error probability is tolerated. Figure 9 shows available computational time as a function of radiation impact rates around values estimated for the two cases. The contour lines show various constant error rates and the horizontal green shaded area ( ms) represents the T1 times of contemporary transmon qubits. In this regime, radiation causes error with a minuscule . In other words, when a calculation is repeated times, one result will be corrupted due to radiation. Most of the modern algorithms running on quantum hardware like variational quantum eigensolver or quantum approximate optimization take less than a milli-second for a single iteration and require a large number of repetitions (). Therefore, we can safely conclude that radiation has negligible effect on modern transmon-based quantum computing platforms even if error-correction fails due to an impact.
V Conclusions
We have performed a comparative study of how radiation may affect transmon qubits hosted at two laboratories with very distinct radiation environments. The regular T1 experiments performed on the same qubit showed no difference in qubit lifetimes of the order of 100 microseconds. The experiments with controlled Thorium sources conducted at the deep underground Gran Sasso laboratory demonstrate the possibility of achieving a good detection efficiency, opening the possibility of using transmons as a particle detector. However, the background rate observed with no active sources at LNGS show higher event rates compared to the simulated values, indicating that other noise sources produce signatures similar to those expected or conjectured to be coming from radiation.
The same protocol when applied to multiple similar qubits in Fermilab (above-ground and with no lead shield) with ambient radiation, shows both increased and reduced rates compared to the response of the maximally shielded qubit at LNGS. These results suggest that these radiation-like events lasting milliseconds are generated by different sources, and that radiation is not a major source of errors in modern transmon qubits. Based on the experimental and simulated results we find that quantum computation will be affected at a maximum of a percent level due to radiation inflicted single-qubit errors when computational time reaches more than 170 ms at fully unshielded facilities and 2.5 s at LNGS. We found no correlation of radiation events to TLS scrambling or qubit excitation.
Future experiments will aim at in-depth understanding of the various error sources which are currently dominating over the radiation induced effects. This will involve further investigations based on fast decay protocols testing different qubit geometries, materials and synchronized measurements across same and different chips. SQMS is focused on enhancing the T1 of transmon qubits and so future measurements will investigate the behavior of qubits with a wide distribution of coherence performance ranging from tens of microseconds to approaching milliseconds. Advancing this knowledge will be critical for the use of superconducting qubits for quantum computation and sensing.
Acknowledgements.
The authors thank the Director and technical staff of the Laboratori Nazionali del Gran Sasso. We are also grateful to the LNGS Computing and Network Service for computing resources and support on U-LITE cluster at LNGS. We are grateful to A. Girardi for the controller of the cryogenic switch, M. Guetti for the help with the whole cryogenic facility (including its continuous upgrades), and M. Iannone for the assembly of the device at LNGS. We are thankful to Theodore C. III and Grzegorz Tatkowski for providing support on experimental preparations and cryogenic operations at FNAL. T.R. thanks Xinyuan You for useful discussions. The work was supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Superconducting Quantum Materials and Systems (SQMS) Center under the contract No. DE-AC02-07CH11359, by the Italian Ministry of Foreign Affairs and International Cooperation, grant number US23GR09, by the Italian Ministry of Research under the PRIN Contract No. 2020h5l338 (“Thin films and radioactivity mitigation to enhance superconducting quantum processors and low temperature particle detectors”) and by the Italian Ministry of Research under the PRIN Contract No. 2022BP4H73 (Chasing Light Dark Matter with Quantum Technologies”).References
- Barends et al. [2014] R. Barends, J. Kelly, A. Megrant, A. Veitia, D. Sank, E. Jeffrey, T. C. White, J. Mutus, A. G. Fowler, B. Campbell, Y. Chen, Z. Chen, B. Chiaro, A. Dunsworth, C. Neill, P. O’Malley, P. Roushan, A. Vainsencher, J. Wenner, A. N. Korotkov, A. N. Cleland, and J. M. Martinis, Superconducting quantum circuits at the surface code threshold for fault tolerance, Nature 508, 500 (2014).
- Arute et al. [2019] F. Arute et al., Quantum supremacy using a programmable superconducting processor, Nature 574, 505 (2019).
- Kjaergaard et al. [2020] M. Kjaergaard, M. E. Schwartz, J. Braumüller, P. Krantz, J. I.-J. Wang, S. Gustavsson, and W. D. Oliver, Superconducting qubits: Current state of play, Annual Review of Condensed Matter Physics 11, 369 (2020).
- [4] Rigetti aspen-m-2 features, https://www.rigetti.com/what-we-build.
- Wu et al. [2021] Y. Wu, W.-S. Bao, S. Cao, F. Chen, M.-C. Chen, X. Chen, T.-H. Chung, H. Deng, Y. Du, D. Fan, M. Gong, C. Guo, C. Guo, S. Guo, L. Han, L. Hong, H.-L. Huang, Y.-H. Huo, L. Li, N. Li, S. Li, Y. Li, F. Liang, C. Lin, J. Lin, H. Qian, D. Qiao, H. Rong, H. Su, L. Sun, L. Wang, S. Wang, D. Wu, Y. Xu, K. Yan, W. Yang, Y. Yang, Y. Ye, J. Yin, C. Ying, J. Yu, C. Zha, C. Zhang, H. Zhang, K. Zhang, Y. Zhang, H. Zhao, Y. Zhao, L. Zhou, Q. Zhu, C.-Y. Lu, C.-Z. Peng, X. Zhu, and J.-W. Pan, Strong quantum computational advantage using a superconducting quantum processor, Phys. Rev. Lett. 127, 180501 (2021).
- Chow et al. [2021] J. Chow et al., Ibm quantum breaks the 100‑qubit processor barrier (2021), .
- Dixit et al. [2021] A. V. Dixit, S. Chakram, K. He, A. Agrawal, R. K. Naik, D. I. Schuster, and A. Chou, Searching for dark matter with a superconducting qubit, Phys. Rev. Lett. 126, 141302 (2021).
- Agrawal et al. [2024] A. Agrawal, A. V. Dixit, T. Roy, S. Chakram, K. He, R. K. Naik, D. I. Schuster, and A. Chou, Stimulated emission of signal photons from dark matter waves, Phys. Rev. Lett. 132, 140801 (2024).
- Day et al. [2003] P. K. Day, H. G. LeDuc, B. A. Mazin, A. Vayonakis, and J. Zmuidzinas, A broadband superconducting detector suitable for use in large arrays, Nature 425, 817 (2003).
- Irwin [1995] K. D. Irwin, An application of electrothermal feedback for high resolution cryogenic particle detection, Applied Physics Letters 66, 1998 (1995).
- Ullom and Bennett [2015] J. N. Ullom and D. A. Bennett, Review of superconducting transition-edge sensors for x-ray and gamma-ray spectroscopy, Superconductor Science and Technology 28, 084003 (2015).
- Charaev et al. [2023] I. Charaev, D. A. Bandurin, A. T. Bollinger, I. Y. Phinney, I. Drozdov, M. Colangelo, B. A. Butters, T. Taniguchi, K. Watanabe, X. He, O. Medeiros, I. Božović, P. Jarillo-Herrero, and K. K. Berggren, Single-photon detection using high-temperature superconductors, Nature Nanotechnology 18, 343 (2023).
- Cardani et al. [2018] L. Cardani, N. Casali, A. Cruciani, H. le Sueur, M. Martinez, F. Bellini, M. Calvo, M. G. Castellano, I. Colantoni, C. Cosmelli, A. D’Addabbo, S. D. Domizio, J. Goupy, L. Minutolo, A. Monfardini, and M. Vignati, Al/ti/al phonon-mediated kids for uv–vis light detection over large areas, Superconductor Science and Technology 31, 075002 (2018).
- Cruciani et al. [2022] A. Cruciani, L. Bandiera, M. Calvo, N. Casali, I. Colantoni, G. Del Castello, M. del Gallo Roccagiovine, D. Delicato, M. Giammei, V. Guidi, J. Goupy, V. Pettinacci, G. Pettinari, M. Romagnoni, M. Tamisari, A. Mazzolari, A. Monfardini, and M. Vignati, BULLKID: Monolithic array of particle absorbers sensed by kinetic inductance detectors, Applied Physics Letters 121, 213504 (2022).
- Mazin et al. [2012] B. A. Mazin, B. Bumble, S. R. Meeker, K. O’Brien, S. McHugh, and E. Langman, A superconducting focal plane array for ultraviolet, optical, and near-infrared astrophysics, Opt. Express 20, 1503 (2012).
- Monfardini et al. [2011] A. Monfardini, A. Benoit, A. Bideaud, L. Swenson, A. Cruciani, P. Camus, C. Hoffmann, F. X. Désert, S. Doyle, P. Ade, P. Mauskopf, C. Tucker, M. Roesch, S. Leclercq, K. F. Schuster, A. Endo, A. Baryshev, J. J. A. Baselmans, L. Ferrari, S. J. C. Yates, O. Bourrion, J. Macias-Perez, C. Vescovi, M. Calvo, and C. Giordano, A dual-band millimeter-wave kinetic inductance camera for the iram 30 m telescope, The Astrophysical Journal Supplement Series 194, 24 (2011).
- Cardani et al. [2020] L. Cardani, N. Casali, G. Catelani, T. Charpentier, M. Clemenza, I. Colantoni, A. Cruciani, L. Gironi, L. Gruenhaupt, D. Gusenkova, et al., Demetra: Suppression of the relaxation induced by radioactivity in superconducting qubits, Journal of Low Temperature Physics 199, 475 (2020).
- McEwen et al. [2024] M. McEwen, K. C. Miao, J. Atalaya, A. Bilmes, A. Crook, J. Bovaird, J. M. Kreikebaum, N. Zobrist, E. Jeffrey, B. Ying, et al., Resisting high-energy impact events through gap engineering in superconducting qubit arrays, arXiv preprint arXiv:2402.15644 (2024).
- Vepsäläinen et al. [2020] A. P. Vepsäläinen et al., Impact of ionizing radiation on superconducting qubit coherence, Nature 584, 551 (2020).
- McEwen et al. [2021] M. McEwen et al., Resolving catastrophic error bursts from cosmic rays in large arrays of superconducting qubits, Nature Physics 18, 107 (2021).
- Li et al. [2024] X.-G. Li, J.-H. Wang, Y.-Y. Jiang, G.-M. Xue, X.-X. Cai, J. Zhou, M. Gong, Z.-F. Liu, S.-Y. Zheng, D.-K. Ma, et al., Direct evidence for cosmic-ray-induced correlated errors in superconducting qubit array, arXiv preprint arXiv:2402.04245 (2024).
- Wilen et al. [2021] C. D. Wilen et al., Correlated charge noise and relaxation errors in superconducting qubits, Nature 594, 369 (2021).
- Thorbeck et al. [2022] T. Thorbeck et al., Tls dynamics in a superconducting qubit due to background ionizing radiation (2022), arXiv:2210.04780 .
- Klimov et al. [2018] P. V. Klimov et al., Fluctuations of energy-relaxation times in superconducting qubits, Phys. Rev. Lett. 121, 090502 (2018).
- de Graaf et al. [2017] S. E. de Graaf et al., Direct identification of dilute surface spins on : Origin of flux noise in quantum circuits, Phys. Rev. Lett. 118, 057703 (2017).
- Burnett et al. [2014] J. Burnett et al., Evidence for interacting two-level systems from the 1/f noise of a superconducting resonator, Nature Communications 5, 4119 (2014).
- Bilmes et al. [2020] A. Bilmes et al., Resolving the positions of defects in superconducting quantum bits, Scientific Reports 10, 3090 (2020).
- Lisenfeld et al. [2019] J. Lisenfeld et al., Electric field spectroscopy of material defects in transmon qubits, npj Quantum Information 5, 105 (2019).
- Martinis et al. [2005] J. M. Martinis et al., Decoherence in josephson qubits from dielectric loss, Phys. Rev. Lett. 95, 210503 (2005).
- Müller et al. [2019] C. Müller et al., Towards understanding two-level-systems in amorphous solids: insights from quantum circuits, Reports on Progress in Physics 82, 124501 (2019).
- McRae et al. [2020] C. R. H. McRae et al., Materials loss measurements using superconducting microwave resonators, Review of Scientific Instruments 91, 091101 (2020).
- Gusenkova et al. [2022] D. Gusenkova et al., Operating in a deep underground facility improves the locking of gradiometric fluxonium qubits at the sweet spots, Applied Physics Letters 120, 054001 (2022).
- Cardani et al. [2021] L. Cardani et al., Reducing the impact of radioactivity on quantum circuits in a deep-underground facility, Nature Communications 12, 10.1038/s41467-021-23032-z (2021).
- Henriques et al. [2019] F. Henriques et al., Phonon traps reduce the quasiparticle density in superconducting circuits, Applied Physics Letters 115, 212601 (2019).
- Martinis [2021] J. M. Martinis, Saving superconducting quantum processors from decay and correlated errors generated by gamma and cosmic rays, npj Quantum Information 7, 90 (2021).
- Iaia et al. [2022] V. Iaia, J. Ku, A. Ballard, C. P. Larson, E. Yelton, C. H. Liu, S. Patel, R. McDermott, and B. L. T. Plourde, Phonon downconversion to suppress correlated errors in superconducting qubits, Nature Communications 13, 6425 (2022).
- Le et al. [2021] D. T. Le et al., Operating a passive on-chip superconducting circulator: Device control and quasiparticle effects, Phys. Rev. Research 3, 043211 (2021).
- Arnquist et al. [2023] I. J. Arnquist, M. L. di Vacri, N. Rocco, R. Saldanha, T. Schlieder, R. Patel, J. Patil, M. Perez, and H. Uka, Ultra-low radioactivity flexible printed cables, EPJ Techniques and Instrumentation 10, 17 (2023).
- Cardani et al. [2023] L. Cardani, I. Colantoni, A. Cruciani, F. De Dominicis, G. D’Imperio, M. Laubenstein, A. Mariani, L. Pagnanini, S. Pirro, C. Tomei, N. Casali, F. Ferroni, D. Frolov, L. Gironi, A. Grassellino, M. Junker, C. Kopas, E. Lachman, C. R. H. McRae, J. Mutus, M. Nastasi, D. P. Pappas, R. Pilipenko, M. Sisti, V. Pettinacci, A. Romanenko, D. Van Zanten, M. Vignati, J. D. Withrow, and N. Z. Zhelev, Disentangling the sources of ionizing radiation in superconducting qubits, The European Physical Journal C 83, 94 (2023).
- Orrell and Loer [2021] J. L. Orrell and B. Loer, Sensor-assisted fault mitigation in quantum computation, Phys. Rev. Applied 16, 024025 (2021).
- Harrington et al. [2024] P. M. Harrington, M. Li, M. Hays, W. V. D. Pontseele, D. Mayer, H. D. Pinckney, F. Contipelli, M. Gingras, B. M. Niedzielski, H. Stickler, J. L. Yoder, M. E. Schwartz, J. A. Grover, K. Serniak, W. D. Oliver, and J. A. Formaggio, Synchronous detection of cosmic rays and correlated errors in superconducting qubit arrays (2024), arXiv:2402.03208 [quant-ph] .
- Mariani et al. [2023] A. Mariani et al., Mitigation of cosmic rays-induced errors in superconducting quantum processors, IEEEQCE23 proceedings (2023).
- Bal et al. [2024] M. Bal, A. A. Murthy, S. Zhu, F. Crisa, X. You, Z. Huang, T. Roy, J. Lee, D. van Zanten, R. Pilipenko, I. Nekrashevich, D. Bafia, Y. Krasnikova, C. J. Kopas, E. O. Lachman, D. Miller, J. Y. Mutus, M. J. Reagor, H. Cansizoglu, J. Marshall, D. P. Pappas, K. Vu, K. Yadavalli, J.-S. Oh, L. Zhou, M. J. Kramer, F. Q. Lecocq, D. P. Goronzy, C. G. Torres-Castanedo, G. Pritchard, V. P. Dravid, J. M. Rondinelli, M. J. Bedzyk, M. C. Hersam, J. Zasadzinski, J. Koch, J. A. Sauls, A. Romanenko, and A. Grassellino, Systematic improvements in transmon qubit coherence enabled by niobium surface encapsulation, npj Quantum Information 10, 43 (2024).
- Yelton et al. [2024] E. Yelton, C. P. Larson, V. Iaia, K. Dodge, G. L. Magna, P. G. Baity, I. V. Pechenezhskiy, R. McDermott, N. Kurinsky, G. Catelani, and B. L. T. Plourde, Modeling phonon-mediated quasiparticle poisoning in superconducting qubit arrays (2024), arXiv:2402.15471 [quant-ph] .
- Martinez et al. [2019] M. Martinez, L. Cardani, N. Casali, A. Cruciani, G. Pettinari, and M. Vignati, Measurements and simulations of athermal phonon transmission from silicon absorbers to aluminum sensors, Phys. Rev. Appl. 11, 064025 (2019).
- Agostinelli et al. [2003] S. Agostinelli et al., GEANT4—a simulation toolkit, Nucl. Instrum. Methods. Phys. Res. A 506, 250 (2003).
- Martinis et al. [2009] J. M. Martinis, M. Ansmann, and J. Aumentado, Energy decay in superconducting josephson-junction qubits from nonequilibrium quasiparticle excitations, Phys. Rev. Lett. 103, 097002 (2009).
- Lenander et al. [2011] M. Lenander, H. Wang, R. C. Bialczak, E. Lucero, M. Mariantoni, M. Neeley, A. D. O’Connell, D. Sank, M. Weides, J. Wenner, T. Yamamoto, Y. Yin, J. Zhao, A. N. Cleland, and J. M. Martinis, Measurement of energy decay in superconducting qubits from nonequilibrium quasiparticles, Phys. Rev. B 84, 024501 (2011).
- Stefanazzi et al. [2022] L. Stefanazzi, K. Treptow, N. Wilcer, C. Stoughton, S. Montella, C. Bradford, G. Cancelo, S. Saxena, H. Arnaldi, S. Sussman, A. Houck, A. Agrawal, H. Zhang, C. Ding, and D. I. Schuster, The qick (quantum instrumentation control kit): Readout and control for qubits and detectors (2022), arXiv:2110.00557 [quant-ph] .
- Fowler et al. [2024] J. W. Fowler, P. Szypryt, R. Bunker, E. R. Edwards, I. F. Florang, J. Gao, A. Giachero, S. F. Hoogerheide, B. Loer, H. P. Mumm, N. Nakamura, G. C. O’Neil, J. L. Orrell, E. M. Scott, J. Stevens, D. S. Swetz, B. A. VanDevender, M. Vissers, and J. N. Ullom, Spectroscopic measurements and models of energy deposition in the substrate of quantum circuits by natural ionizing radiation (2024), arXiv:2404.10866 .
- Loer et al. [2024] B. Loer, P. M. Harrington, B. Archambault, E. Fuller, B. Pierson, I. Arnquist, K. Harouaka, T. D. Schlieder, D. K. Kim, A. J. Melville, B. M. Niedzielski, J. K. Yoder, K. Serniak, W. D. Oliver, J. L. Orrell, R. Bunker, B. A. VanDevender, and M. Warner, Abatement of ionizing radiation for superconducting quantum devices (2024), arXiv:2403.01032 .
- Linehan et al. [2024] R. Linehan, I. Hernandez, D. J. Temples, S. Q. Dang, D. Baxter, L. Hsu, E. Figueroa-Feliciano, R. Khatiwada, K. Anyang, D. Bowring, G. Bratrud, G. Cancelo, A. Chou, R. Gualtieri, K. Stifter, and S. Sussman, Estimating the energy threshold of phonon-mediated superconducting qubit detectors operated in an energy-relaxation sensing scheme (2024), arXiv:2404.04423 .
Author Contributions
M.B. and F.C. designed and fabricated the transmon devices. S.Z. performed initial characterization of the qubits. A.C. T.R., F.D.D. and D.H. run the experiments. N.C. developed the Bernoulli algorithm. A.M., F.D.D, T.R. and L.C. performed the data analysis. D.V.Z. helped in writing acquisition software. A.M. performed GEANT4 simulations. I.C., A.P., L.P. and S.P. oversaw the cryostat upgrade for the measurements and the cryogenic operations. T.R., A.M., F.D.D., A.G. and L.C. wrote the manuscript with inputs from all the co-authors in interpreting the results. T.R., A.G. and L.C. conceived and supervised the whole project.
SUPPLEMENT
V.1 Device Fabrication
The devices are fabricated on a 432 m thick -plane HEMEX grade double-polished sapphire substrate with 2-inch diameter. The substrate preparation, base-layer (Nb) deposition and Josephson junction evaporation are similar to the fabrication methods followed in Ref. 43. In order to prevent oxidation of the Niobium, a capping layer of thickness nm is deposited using the same sputtering tool without breaking the vacuum. Gold and Tantalum are used as capping layers for devices Q1 and Q2-Q4 respectively (on two different chips). Device Q1 is first measured at FNAL and then sent to LNGS. The other qubit samples (Q2-Q4) are measured at FNAL. In Table S1, we list the measured parameters of the devices and Fig. S1 shows the relative position of those.
Parameter | Q1 | Q2 | Q3 | Q4 | Units |
Material | Nb/Au | Nb/Ta | Nb/Ta | Nb/Ta | N/A |
Qubit frequency | 4717.4 | 4455.4 | 4451.3 | 4294.8 | MHz |
Readout frequency | 7206.8 | 7055.0 | 6886.5 | 6714.5 | MHz |
Qubit pulse length | 0.150 | 0.091 | 0.124 | 0.160 | s |
Qubit average T1 | 84 | 141 | 131 | 214 | s |
Readout pulse length | 4.5 | 3.8 | 4.0 | 8.0 | s |
Waiting period | 5.0 | 10.0 | 5.0 | 5.0 | s |
Cooldown period | 50.0 | 70.0 | 70.0 | 10.0 | s |
One iteration period | 64.550 | 87.929 | 84.324 | 31.660 | s |
V.2 Measurement Setup
Our measurement setups at the two locations are shown in Fig. S2. The transmon qubit is measured in the hangar geometry, where a CPW readout resonator is coupled to a transmission line. The pulses for both qubit control and readout are sent through a common input line. The input line is attenuated by 60-70 dB distributed between the 4-K and mixing chamber stages. Copper-body 20 dB attenuators are used at the base plate plate for better thermalization. At LNGS, the input line is also filtered through a 8 GHz reflective low-pass filter and a strong infrared filter (QMC-CRYOIRF-002) to reduce high-frequency noise. The superconducting output line is protected from the incoming noise through a 4-8 GHz triple junctions isolator, an 8 GHz low-pass filter and a weak infrared filter (QMC-CRYOIRF-001). At FNAL, both the input and output lines are filtered through 11 GHz reflective low-pass filters. The superconducting output line also has two 4-8 GHz isolators. The device boxes at both locations are protected from stray magnetic fields using cryoperm shields. The output signal is first amplified through a commercial HEMT amplifier at the 4-K plate, followed by a room-temperature amplifier outside the fridge before digitization. The qubit control and readout pulses are generated digitally using a RFSoC ZCU 216 evaluation board and the QICK [49] software.
V.3 Simulation of Interaction Rates
We use the GEANT4-based Monte Carlo simulation developed in Cardani et al. [39] within the SQMS Round Robin project with updated parameters. The aim is to predict both the rate of impacts and the amount of energy released in the Sapphire substrate. The geometry implemented in the simulation, shown in Fig. S3, includes (from outside to inside):
a.
An 80 80 cm2, 5 cm-thick lead base;
b.
An 80 80 cm2, 5 cm-thick copper base;
c.
1/2 round of 10 20 cm2, 5 cm-thick Pb bricks;
d.
1 round of 10 20 cm2, 5 cm-thick Pb bricks;
e.
1 round of 10 66.5 cm2, 2 cm-thick Cu bars;
f.
A two layers Permimphy magnetic shielding (Outer layer: 55.6 cm (OD) 91.3 cm, 1.5 mm-thick; Inner layer: 51 cm (ID) 89 cm, 1.5 mm-thick);
g.
A 48.1 cm (OD) 119.5 cm simplified version of the cryostat with internal vessels;
h.
A CryoPerm® magnetic shield below the mixing chamber;
j.
A 3 cm-thick lead plate at 10 mK, placed between the sample and the electronic components used as internal gammas shield;
k.
A 30.23 30.23 6.3 mm3 gold-plated copper box hosting the chip.
The chip substrate (7.5 7.5 mm2, 432 m-thick HEMEX Sapphire) is considered the only active volume, where we store tracks and record the energy released by each simulated interaction. The main external contributions, namely environmental gammas and muons, are generated according to the methods described in our previous work [39]. For each source, and for both FNAL and INFN-LNGS laboratories, we estimate the rate of interactions in the chip by scaling the number of recorded events to the flux measurements reported in Section 3 of Ref. [39].
Regarding close sources of radioactivity, we simulate the radioactive decays of relevant isotopes (40K, 232Th and 238U chains) uniformly distributed within the volume of the two Printed Circuit Boards (RO4003C produced by ROGERS Corporation) located at the opposite ends of the qubit chip. PCBs indeed constitute the only significant contribution from close sources, as detailed in [39]. We then estimate the rate of interactions in the chip by scaling the number of recorded events to the activities reported in Table S2. These activities are determined by converting the concentrations measured using Inductively Coupled Plasma Mass Spectrometry (ICP-MS) at the Chemistry facility in INFN-LNGS.
Source | 40K | 232Th | 238U |
mBq/kg | mBq/kg | mBq/kg | |
PCB | (518.8 259.4) | (164.0 41.0) | (161.2 49.6) |
Figure S4 presents the results of the Monte Carlo simulation, showing the rate of events as a function of the energy deposited in the qubit substrate for both FNAL and INFN-LNGS facilities. The event rates for each radiation source are reported in Table 1. Notably, the overall rate is expected to be approximately an order of magnitude lower at INFN-LNGS (4 @INFN-lNGS vs. 57 @FNAL), where the dominant contribution is from the PCBs.
To investigate the rate in the presence of the Thorium radioactive sources used at INFN-LNGS, we generate radioactive decays from a point-like 232Th radioactive source positioned between the external magnetic shield and the copper bars, approximately at the same height of the chip. For this simulation, we consider only gamma emitters in the chain (228Ra, 228Ac, 212Pb, 212Bi and 208Tl), since and particles cannot reach the chip as they are shielded by the cryostat. As in the previous case, we determine the rate of interactions in the chip by scaling the number of recorded events to the activities reported in Table S3, which correspond to the Thorium source activities employed in the measurements. The rate of events in the chip, plotted as a function of the deposited energy for the various activities, is illustrated in Fig. S3.
Activity | Rate |
kBq | events/sec |
44.2 | 0.12 0.01 |
75.9 | 0.20 0.02 |
125.4 | 0.34 0.03 |
161.0 | 0.43 0.04 |
V.4 Qubit State Discrimination
The I/Q values (in-phase and quadrature-phase signals) obtained for each measurement depend on the state of the qubit. When a large counts of measurements are plotted on the I-Q plane, we usually observe two blobs corresponding to and . However, we sometimes observe signals corresponding the transmon being in the second excited state as shown in Fig. S6 and we discard those measurements in the analysis.
The blobs corresponding to and are rotated in order to have the centers at the same value of Q and projected onto the I axis. In this way it is possible to discriminate between the two states by putting a threshold value only on one (I) quadrature.
The populations of the states have been estimated by doing a simultaneous fit on the two quadratures with the RooFit toolkit. The fit function used was a sum of 2D gaussians, one for each blob we were observing. The data was divided into traces of approx. 1 minute duration, and for each of them the simultaneous fit was performed. In this way we were able to identify periods in which the qubit was showing a high population of the state and reject them.
V.5 Data Analysis
In our analysis we selected events where the probability of observing a certain number of zeros and ones, calculated according to the Binomial distribution, was below a predefined threshold. Specifically, if the probability of observing such a series of zeros and ones fell below 1%, corresponding to an absolute logarithm () greater than 4.6, the software trigger was fired, indicating the presence of a potential signal. Once triggered, these data wedre recorded for further analysis.
By selecting pulses with , we observe a trigger rate of 5–6 events/second in the runs performed with the radioactive sources, significantly higher than the expected rate from radioactivity (0.1–0.4 events per second depending on the source activity), suggesting that the majority of triggers are due to noise. To address this, we computed additional variables to refine our data selection. Specifically, we calculated:
1.
The value and RMS of the baseline before the pulse;
2.
The maximum value of ;
3.
The pulse length, defined as the number of points above threshold;
4.
The position of the maximum.
The baseline and its RMS were computed using the first 100 points preceding the trigger. We then removed pulses occurring on the tail of a preceding pulse or those with a “noisy” baseline (see Fig. 4) by requiring a baseline value lower or equal to its mean value + 3. Additionally, we ensured that the position of the maximum was close to the trigger to avoid re-triggering of the same pulse. Since radioactivity typically results in long pulses (at least 1 ms), we further refined our selection by choosing pulses with length greater than 50 points, corresponding to approximately 3.5 ms under our data-taking conditions.
Finally, we conducted a study to assess the suitability of the window size used for computing the Binomial probability. Our initial choice of 50 points stemmed from the analysis of typical events detected in calibration runs. By averaging tens of these events, we determined that the qubit returned to equilibrium with a characteristic time constant () of 0.78 ms. To ensure complete recovery of the qubit, we adopted a window size corresponding to 5, or approximately 4 ms, which translates to 56 points given our sampling frequency of about 0.07 ms-1. We subsequently performed the analysis using window sizes of 20, 30, 70, and 90 points. Our findings confirmed that, for all these window sizes, the detection efficiency could not reach values above 70. Indeed, when the window size is too small, only a fraction of the signal can be captured within the window. Conversely, if the window size exceeds the signal length, the number of zeros associated with random qubit de-excitation becomes significant, thus deteriorating the signal-to-noise ratio.

### v2

@@
 Previous studies have focused on understanding possible correlated effects over time and distance due to cosmic rays.
-In this study, for the first time, we directly compare the response of a transmon qubit measured initially at the Fermilab SQMS above-ground facilities and then at the deep underground Gran Sasso Laboratory (INFN-LNGS, Italy).
-We observe same average qubit lifetime T1 of roughly 80 microseconds at above and underground facilities.
-We then apply a fast decay detection protocol and investigate the time structure, sensitivity and relative rates of triggered events due to radiation versus intrinsic noise, comparing above and underground performance of several high-coherence qubits.
-Using gamma sources of variable activity we calibrate the response of the qubit to different levels of radiation in an environment with minimal background radiation.
-Results indicate that qubits respond to a strong gamma source and it is possible to detect particle impacts.
-However, when comparing above and underground results, we do not observe a difference in radiation induced-like events for these sapphire and niobium-based transmon qubits.
-We conclude that the majority of these events are not radiation related and to be attributed to other noise sources which by far dominate single qubit errors in modern transmon qubits.
-In recent decades, superconducting circuits have emerged as a leading technology explored in the fields of quantum information [1, 2, 3, 4, 5, 6], photon and particle detectors [7, 8, 9, 10, 11, 12, 13, 14, 15, 16].
+In this study, for the first time, we directly compare the response of a transmon qubit measured initially at the Fermilab SQMS above-ground facilities and then at the deep-underground Gran Sasso Laboratory (INFN-LNGS, Italy).
+We observe the same average qubit lifetime T1 of roughly 80 microseconds at both above and underground sites despite having very different radiation levels.
+We then apply a fast decay detection protocol and investigate the time structure, sensitivity and relative rates of triggered events due to radiation versus intrinsic noise, comparing above and underground performance of several high-coherence qubits with similar design.
+When comparing above and underground results, we do not observe a significant difference in radiation induced-like events for these sapphire and niobium-based transmon qubits.
+We conclude that the majority of these events are not radiation-related and can be attributed to other noise sources, which predominantly contribute to single qubit errors in modern transmon qubits.
+Using gamma sources with increasing activity, we assess the response of the qubit to different levels of radiation in an environment with minimal background radiation.
+Results indicate that qubits respond to gamma sources, demonstrating the possibility to detect particle impacts.
+However, radiation events can be observed only by significantly reducing other noise sources, which compromises detection efficiency.
+In recent decades, superconducting circuits have emerged as a leading technology for applications in the fields of quantum information [1, 2, 3, 4, 5, 6], photon, and particle detectors [7, 8, 9, 10, 11, 12, 13, 14, 15, 16].
 In the past years, several research groups have started investigating the effect of particle interactions on the performance of quantum bits (qubits) [17, 18].
-There are many studies investigating these potential effects in superconducting quantum devices of various geometries and materials.
-Some early work pointed to the lifetime of a transmon qubit potentially decreasing when exposed to a strong radioactive source [19].
+There are many studies examining these potential effects in superconducting quantum devices of various geometries and materials.
+Some early work indicated that the lifetime of a transmon qubit could potentially decrease when exposed to a strong radioactive source [19].
 Moreover, the effect of high energy deposits can rapidly propagate from the initial location of deposition, affecting multiple qubits and potentially leading to a general chip failure [20, 21].
 C.D.
-Wilen et al.[22] observed that the rate of charge jumps experienced by an array of qubits (operated as electrometers) was consistent with the rate expected from interactions due to ambient radioactivity and cosmic rays.
+Wilen et al.
+[22] observed that the rate of charge jumps experienced by an array of charge-sensitive qubits (operated as electrometers) was consistent with the rate expected from interactions due to ambient radioactivity and cosmic rays.
 Thorbeck et al.
-[23] did not observe significant radiation effects, but concluded that the impacts due to radioactivity may induce frequency jumps in a dominant loss mechanism, namely two-level systems (TLS) [24, 25, 26, 27, 28, 29, 30, 31], possibly affecting qubit stability.
+[23] did not observe significant radiation effects, but concluded that the impacts due to ionizing radiation may induce frequency jumps in a dominant loss mechanism, namely two-level systems (TLSs) [24, 25, 26, 27, 28, 29, 30, 31], possibly affecting qubit stability.
 Other studies have shown that suppressing radioactivity can enhance the performance of superconducting circuits.
-Operating the device in a radio-pure facility has allowed to significantly improve the locking of a fluxonium qubit at the sweet spot [32] and enhanced the internal quality factor of superconducting resonators [33].
-The latter publication also enabled the determination that the substrate on which qubits are deposited, rather than the qubits themselves, is the primary target for particle interactions.
+Operating the device in a radio-pure facility has allowed for the suppression of correlated errors [32] significantly improving the locking of a fluxonium qubit at the sweet spot [33] and enhancing the internal quality factor of superconducting resonators [34].
+The latter publication also determined that the substrate on which qubits are deposited, rather than the qubits themselves, is the primary target for particle interactions.
 These findings have motivated efforts in further directions aimed at studying and mitigating possible effects of radioactivity.
-Promising results have been achieved by equipping the chip with “traps” designed to absorb phonons produced (also) by radioactive interactions [34, 35, 36] and using gap engineering [18].
+Promising results have been achieved by equipping the chip with “traps” designed to absorb phonons produced (also) by radioactive interactions [35, 36, 37] and using gap engineering [18, 38].
 Some groups are also focusing on the components used to construct and operate qubits [39, 40], as it has been demonstrated that some of them, such as the Printed Circuit Boards (PCBs), can produce a sizeable rate of events in the qubit chip [41].
-Furthermore, it is also worth mentioning recent proposals of equipping qubits with a device specialised in recognizing particle interactions [40] or with external muon detectors [41, 21, 42].
-In this work, for the first time, we study charge-insensitive transmon qubit devices in two laboratories with very distinct radiation environments, as a step of the Superconducting Quantum Materials and Systems (SQMS) Center Round Robin project [39], which aims at studying identical qubits measured at multiple test sites around the world.
-We first compare average relaxation times measured over several hours, and then employ a fast decay detection protocol to identify reductions in relaxation time at rates faster than 10 kHz and analyze time-bin correlations to determine the characteristics of the events.
-We study the response of the qubit in a controlled environment, by exposing it to gamma-ray sources of increasing activity.
-Using the same protocol and analysis strategy, we investigate similar qubits hosted above ground in a typical laboratory environment, to compare effect of radiation on transmon qubits performance and behavior, and estimate timescales at which radiation may become concerning for quantum computational errors.
+Furthermore, it is also worth mentioning recent proposals of equipping qubits with a device specialized in recognizing particle interactions [42] or with external muon detectors [38, 21, 43].
+In this work, for the first time, we study charge-insensitive transmon qubit devices in two laboratories with very distinct radiation environments, as part of the Superconducting Quantum Materials and Systems (SQMS) Center Round Robin project [41], which aims to study identical qubits measured at multiple test sites around the world.
+We first compare the average relaxation time of a qubit measured over several hours both above ground at Fermilab and deep underground at Gran Sasso Laboratory.
+We then employ a fast decay detection protocol to investigate radiation-induced events in similar transmon qubits and analyze the time structure, sensitivity, and relative rates of triggered events due to radiation versus intrinsic noise.
+To further study the qubit’s response to radiation, we expose the superconducting chip to gamma-ray sources with increasing activity in a controlled environment with minimal background radiation and estimate timescales at which radiation may become concerning for quantum computational errors.
 I Experimental locations
-To perform the comparative study, we operated similar transmon chips (charge-insensitive) in two different laboratories: the Quantum Garage of the SQMS Center at Fermilab (FNAL, Illinois, US), and the Ieti Facility located in the deep underground Gran Sasso Laboratory of INFN (INFN-LNGS, L’Aquila, Italy).
+To perform the comparative study, we operated similar transmon chips (charge-insensitive) in two different laboratories: the Quantum Garage of the SQMS Center at Fermilab (FNAL, Illinois, US), and the Ieti Facility located in the deep-underground Gran Sasso Laboratory of INFN (INFN-LNGS, L’Aquila, Italy).
 The chip consisted of a 432 m thick, 7.57.5 mm2 wide HEMEX grade Sapphire substrate.
+It is noteworthy that the entire substrate is covered in Niobium creating a ground plane; this may act as a phonon trap, diminishing the phonon propagation and absorption in the active part of the qubit [44, 45].
 Eight Niobium transmons with different geometries were deposited on the chip (see Supplementary section A); their surface was capped using a 10 nm thick layer of Gold (measured at both sites) or Tantalum (measured at FNAL) to reduce losses due to the formation of Nb2O5 [46].
-The capping technique allowed to achieve excellent T1 of many tens and up to hundreds of microseconds, which is important to study the change in qubit’s lifetime at microsecond timescales.
-We characterized several qubits of these chips operated at FNAL and at INFN-LNGS, and performed comparisons for either an identical qubit or qubits of similar geometries and materials employed.
-The detailed parameters of these devices are provided in the Supplementary Section A.
-Additionally, it is noteworthy that the entire substrate is covered in Niobium creating a ground plane; this may act as a phonon trap, diminishing the phonon propagation and absorption in the active part of the qubit [44, 45].
+The capping technique allowed us to achieve excellent T1 of many tens and up to hundreds of microseconds, which is crucial for studying changes in a qubit’s lifetime at microsecond timescales.
+We characterized several qubits of these chips operated at FNAL and at INFN-LNGS, and performed comparisons for either an identical qubit (standard T1 experiment only) or qubits of similar geometries and materials.
 To predict the rate of impacts due to ambient radiation, we conducted a Monte Carlo simulation using the GEANT4 framework [47] developed in Ref.
 [41].
-We anticipate the total rate of radioactive events interacting with the chip at FNAL and INFN-LNGS to be events/sec and events/sec, respectively.
-Table 1 provides a summary of the simulation results, with detailed information provided in the Supplementary Section C.
+We anticipate the total rate of radiation events interacting within the chip at FNAL and INFN-LNGS to be events/sec and events/sec, respectively.
+Table 1 provides a summary of the simulation results, with detailed information provided in the Supplementary Section B.
 Source | FNAL | INFN-LNGS |
@@
 At FNAL (see Fig.
-1(a)), the rate is primarily dominated by rays produced by the decay of naturally radioactive isotopes present in the laboratory environment.
+1(a), left panel), the rate is primarily dominated by rays produced by the decay of naturally radioactive isotopes present in the laboratory environment.
 However, muons generated by cosmic rays also contribute significantly to the overall rate.
 In contrast, at INFN-LNGS (see Fig.
-1(a)), muons are suppressed by six orders of magnitude thanks to the 1.4 km thick rock overburden.
+1(a), right panel), muons are suppressed by six orders of magnitude thanks to the 1.4 km thick rock overburden.
 Moreover, the radioactivity of the laboratory is substantially reduced by copper and lead shields positioned both inside and around the refrigerator.
@@
 While this rate is negligible at FNAL, it becomes the dominant source of radioactivity at INFN-LNGS, where all external sources of radioactivity are suppressed.
-Furthermore, at INFN-LNGS, we utilize Thorium radioactive sources to increase the event rate up to one event every two seconds and investigate the chip behaviour in a “high”-radiation environment.
+Furthermore, at INFN-LNGS, we utilize calibrated Thorium radioactive sources to increase the event rate up to one event every two seconds and investigate the chip behaviour in a “high”-radiation environment.
 II Effect of radiation impact
-When a radioactive particle impinges on the qubit chip, it releases energy.
+When a particle impinges on the qubit chip, it releases energy.
 Muons create long tracks across the chip, whereas rays interact through photo-electric absorption or (mainly via) Compton scattering, producing short-track electrons.
@@
 Since the metallic pads of the transmon host a large number of quasiparticles that take some time to dissipate through normal metal (PCB leads) or lower-energy phonon scattering, these tunneling events can continue to occur, causing the qubit to suffer from strong relaxation for some time.
-Previous studies claim this low T1 periods to last from one to several tens of milli-seconds [21, 20].
+Previous studies claim this low T1 periods last from one to several tens of milliseconds [21, 20].
 We investigated these type of events in our studies.
-We first performed standard T1 experiments on both underground and above ground locations on the identical qubit Q1 and monitored the relaxation times for close to an hour.
-These experiments were slow and each data point took about 50 seconds to complete.
+We first performed standard T1 experiments on both underground and above-ground locations on the identical qubit Q1 and monitored the relaxation times for close to an hour.
+These experiments were slow, and each data point took about 50 seconds to complete.
 The results, plotted in Fig.
-1(b), show similar mean values and fluctuations, typical for transmon qubits [43], demonstrating that radiation has no significant effect at current state-of-the-art transmon qubit lifetimes of 100 microseconds.
+1(b), show similar mean values and fluctuations, typical for transmon qubits [46], demonstrating that radiation has no significant effect at current standard transmon qubit lifetimes of 100 microseconds.
 No sudden drop in T1 values was observed.
-This was expected as the standard T1 experiments, due to its averaging nature, are not sensitive to fluctuations at milli-second timescales.
+This was expected as the standard T1 experiments, due to their averaging nature, are not sensitive to fluctuations at millisecond timescales.
 In order to study possible events at such a time resolution we employed a protocol, tailored for single qubits, utilizing fast decay measurements [20, 18].
-Datasets | acquisition time | live-time |
-[min] | [min] | |
-Shielded chip | 676.0 | 28.2 |
-Shielded chip (2) | – | 124.0 |
-44 kBq Th Source | 619.5 | 114.0 |
-76 kBq Th Source | 111.7 | 46.3 |
-125 kBq Th Source | 60.9 | 28.2 |
-161 kBq Th Source | 60.9 | 54.3 |
-FNAL Q2 | 234.5 | 104.0 |
-FNAL Q3 | 449.7 | 141.9 |
-FNAL Q4 | 242.7 | 25.9 |
 The fast detection protocol illustrated in Fig.
 2(b) begins by resetting the qubit in its excited state .
-This is accomplished by initially measuring the qubit and applying a conditional -pulse if the qubit is found to be in state via fast feedback.
+This is accomplished by initially measuring the qubit and applying a conditional -pulse if the qubit is found to be in the state via fast feedback.
 Following a waiting period lasting a few microseconds, the qubit is measured again to determine whether it is in state or .
@@
 Theoretically, the qubit relaxes to state only a small fraction of the time, given by .
-However, a radiation event significantly reduces the T1, making it highly probable for the qubit to be detected in state .
+However, a radiation event significantly reduces T1, making it highly probable for the qubit to be detected in state .
 The sequence is then repeated after another short period, allowing the chip to cool down.
@@
 2(c).
-In our measurements, we kept the waiting period fixed to 5 s while other parameters like -pulse and readout pulse lengths were similar and reported in the Supplementary Section A.
-The cooldown period is chosen to ensure that the transmon does not exhibit excessive population in the second excited state .
+In our measurements, we select a fixed waiting period of 5 or 10 s; the readout pulse has a length ranging from 3.8 to 8 s (the individual parameters of each measurement are reported in Table S1 of the Supplementary Section A).
+The -pulse has a negligible length ( 1 s).
+The cool-down period (10 to 70 s) is chosen to ensure that the transmon does not exhibit excessive population in the second excited state .
 Nevertheless, a simultaneous fit of the , and the second excited state showed that there were many moments in which the (and sometimes even the third excited state ) states were populated.
-To ensure data consistency, we excluded these time intervals.
+To ensure data consistency, we exclude these time intervals.
 A thorough investigation of this phenomenon will be mandatory for applications in sensing, as this selection significantly limits the measurement’s live-time (see Table 2).
-The readout scheme employed is detailed in the Supplementary Section B.
+Datasets | acquisition time | live-time |
+[min] | [min] | |
+Shielded chip | 736.7 | 28.2 |
+44 kBq Th Source | 620.7 | 114.0 |
+76 kBq Th Source | 124.1 | 46.3 |
+125 kBq Th Source | 62.1 | 23.7 |
+161 kBq Th Source | 62.1 | 62.1 |
+FNAL Q2 | 234.6 | 104.1 |
+FNAL Q3 | 449.7 | 92.8 |
+FNAL Q4 | 243.1 | 25.9 |
+The readout scheme employed is detailed in the Supplementary Section C.
 The input signal, filtered through attenuators, low-pass filters, and infrared filters, is sent to the qubit; the output signal is fed into cryogenic triple-junction isolators, and then amplified via a cryogenic HEMT followed by a room temperature amplifier.
@@
 We acquired the data for several hours containing several hundred million readouts.
-During the initial analysis, we simply computed the probability of finding a qubit in for a fixed time window and plot as we moved the window (see Fig.
+During the initial analysis, we simply computed the probability of finding a qubit in for a fixed time window and plotted as we moved the window (see Fig.
 2(c)).
@@
 3(b).
-These pulses generally have a sharp rise but the fall times vary between different instances.
+These pulses generally have a sharp rise, but the fall times vary between different instances.
 As depicted in Fig.
-1(b), we found pulses of similar time-profile at both locations irrespective of the specific qubit, pointing to the fact that the drop in relaxation time is not tied to a specific transmon.
-All pulses observed in our studies last less than ten milli-seconds in agreement with the study in Ref.
-21 and thus these pulses require multiple qubit measurements within a milli-second for being detected.
+1(b), we found pulses of similar time-profile at both locations irrespective of the specific qubit, indicating that the drop in relaxation time is not tied to a specific transmon.
+All pulses observed in our studies last less than ten milliseconds in agreement with the study in Ref.
+21, and thus these pulses require multiple qubit measurements within a millisecond for being detected.
 For example, the pulses in Fig.
-1(c) are visible when a one ms time-window is selected for computing and the signal vanishes when a course time-window of 100 ms is chosen.
+1(c) are visible when a one ms time-window is selected for computing and the signal vanishes when a coarse time-window of 100 ms is chosen.
 Clearly, a slow measurement scheme will not be able to resolve individual radiation impact events.
-In other words, standard T1 measurements, due to its slow nature, will not be affected by radiation impacts (happening at intervals of several seconds) as demonstrated earlier.
+In other words, standard T1 measurements, due to their slow nature, will not be affected by radiation impacts (happening at intervals of several seconds).
 A simple detection strategy based on triggering pulses on the data by setting a constant threshold, however, suffers from the problem that a large threshold misses many events, whereas, a lower value picks up a lot of noise arising inherently from the qubit measurement.
-Furthermore, such a noise is not stable in time leading to unexpected variations when setting a simple constant threshold.
+Furthermore, such a noise is not stable in time leading to unexpected variations in the rate.
 As shown in Fig.
@@
 These periods of instability can last from sub-seconds to a minute and will affect regular T1 experiments.
-Even though it appears to be related to TLS scrambling, we currently do not have a clear understanding of the source of this behavior and it will be a subject of future studies.
-Consequently, to improve the detection efficiency, we develop a modified strategy discussed below.
-III Pulse Reconstruction
-For the event detection, we developed an approach based on the probability of observing a certain stream of and .
-Initially, we performed a simultaneous Gaussian fit to the distribution of I/Q of , , and to correct for time instabilities and removed traces in which the qubit turned out to be in the second excited state .
-The clouds corresponding to and are then rotated in order to be aligned along one quadrature and projected onto the other.
+Even though it appears to be related to TLS scrambling, we currently do not have a clear understanding of the source of this behavior, and it will be a subject of future studies.
+To better identify radiation-induced events, we developed a modified strategy discussed below.
+III Event Selection
+Several considerations are applied to the analysis for the selection of events.
+Initially, we perform a simultaneous 2D Gaussian fit to the distribution of I/Q values for the states , , and .
+To correct for time instabilities, we remove traces where the qubit is found in the second excited state .
+The clouds corresponding to and are then rotated to align along one quadrature and projected onto the other.
 Further information can be found in the Supplementary Section D.
-We then performed a Gaussian fit to determine the point at which and were maximally separated, enabling us to convert the measured I/Q values to a sequence of 0 (for ) and 1 (for ).
-This stream of zeros and ones was further processed to construct pulses for event detection.
-We started with the assumption that, for each measurement, the probability of the qubit remaining in state is equal to .
-Since, our readouts took several micro-seconds, we added that to the waiting period to accurately compute .
-Moreover, we found that the qubit relaxations varied by 25-30 over the data acquisition period (see Fig.
-1(b)), consistent with the typical behavior of transmon devices [43].
-Therefore, was computed using points (corresponding to approximately one minute worth of data) and used to estimate the probability for the qubit to be in state .
-We began by picking the first 50 points of the trace and computing the probability for observing the measured counts of zeros and ones according to the Binomial probability distribution.
-Subsequently, we moved to the next point of the trace and repeated the calculation.
-Given that the probability for the qubit to be in state was approximately 85–90%, the probability of observing a large number of zeros in a subsequent series of measurements was very small.
-Finally, we computed the absolute logarithm () of this probability, simply to avoid working with very small numbers.
-If the probability of observing this series of events was smaller than 1 (), the software trigger was fired, and we saved a pulse with a length of 450 points, 100 of which were acquired before the pulse to monitor the status of the qubit before the interaction.
-Typical triggered pulses are shown in Fig.
+Once the I/Q clouds are projected onto one quadrature, a threshold is chosen in order to convert the measured I/Q values into a sequence of 0s (for ) and 1s (for ).
+Since the two clouds are not completely separated, every value of the threshold will lead to some errors in the state identification.
+In our analysis, we calculate the fraction of measurements for each state falling on the wrong side of the threshold (misidentification) and choose the threshold value that minimizes the sum of these two fractions (total error).
+This is done by fitting the clouds with 1D Gaussian functions and integrating the tails falling on the wrong side of the threshold, varying the candidate threshold value until the minimum is found.
+The resulting threshold values allow for the correct identification of the qubit state in 90% to 95% of measurements, with fluctuations of a few percentage points depending on the specific qubit and dataset considered.
+The sequence of 0s and 1s is then organized into traces of points.
+We expect that radiation-induced events last at least one millisecond, which, given the sampling periods reported in Table 3, corresponds to approximately 12–35 zeros.
+According to the same table, we anticipate a baseline of 8–15% of zeros arising from spontaneous qubit decay.
+Therefore, we require the detection of at least four consecutive zeros to trigger an event.
+This criterion ensures reasonable signal efficiency while significantly reducing false triggers caused by qubit spontaneous decay and other types of noise.
+Indeed, the number of times we measure the qubit in state can be due to different phenomena, which can be summarized as follows:
+1.
+Qubit spontaneous decay: The probability of finding the qubit in state can be roughly estimated by , where is an “effective” relaxation time computed using points (corresponding to approximately one minute of data) to monitor changes over time, and is the waiting period (see Table 3) between the state preparation in and end of readout.
+We then use the average of the values obtained from each data segment to estimate the probability of the qubit being in state for each dataset.
+Further details can be found in Supplementary Section E.
+2.
+Ionizing radiation: We use Geant4-based Monte Carlo simulations to predict the number of impacts on the qubit chip.
+3.
+Other sources of noise: These involve electronic noise, luminescence or relaxation of materials at low temperature [51] (known sources of quasiparticle poisoning include the holder or the interface between the holder and the chip [52, 53, 54] and relaxations of metal films [55, 53]), vibrations [56], etc.
+Datasets | TS | r | |||
+[s] | [s] | [%] | [s] | [events/sec] | |
+Shielded chip | 14 | 118.6 | 11.1 | 73.7 | 2.3 0.4 |
+44 kBq Source | 11 | 76.3 | 13.4 | 67.7 | 5.4 1.1 |
+76 kBq Source | 11 | 79.9 | 12.9 | 67.7 | 4.6 0.9 |
+125 kBq Source | 11 | 68.9 | 14.8 | 67.7 | 7.2 1.5 |
+161 kBq Source | 11 | 74.1 | 13.8 | 67.7 | 6.1 1.2 |
+FNAL Q2 | 14 | 162.7 | 8.2 | 88.0 | 0.5 0.1 |
+FNAL Q3 | 9 | 63.8 | 13.2 | 84.3 | 4.1 0.8 |
+FNAL Q4 | 13 | 165.5 | 7.6 | 31.7 | 1.1 0.2 |
+Once the trigger is fired, we acquire a 160-point window divided as follows:
+A 60-point long () pre-trigger region to estimate the average number of zeros (noise) before the events;
+A 40-point long () dead region to avoid signal leakage into the pre-trigger region;
+A 60-point long () signal region. Specifically, we acquire 10 points before the trigger and 50 after the trigger.
+The choice of the window length was based on the signal features.
+Given a sampling period of 0.03–0.09 milliseconds, and considering that an event lasts a few milliseconds (see Fig.
+1), we expect the event to be fully contained within a 60-point window.
+To prevent re-triggering, we apply a dead-time of 51 points after the trigger.
+As shown in Table 3, we expect from 0.5 up to 7 events per second due to qubit spontaneous decay alone, depending on the dataset.
+These rates are about one to two orders of magnitude higher than those expected from ambient radiation (0.004 and 0.057 events/sec at LNGS and FNAL, respectively) or gamma ray sources (0.12 to 0.43 events/sec).
+To identify events caused by ionizing radiation, we employ a systematic approach aimed at minimizing contributions from qubit spontaneous decay.
+For each dataset, we compute the probability of observing more than a certain number of zeros in a sequence of 60 measurements, corresponding to the length of the signal window.
+This probability is derived from the cumulative distribution function (CDF) of the Binomial distribution calculated over 60 points.
+Based on these probabilities, we set thresholds on the minimum number of zeros required within the signal region to significantly reduce the expected rate of qubit spontaneous decay, targeting values approximately an order of magnitude lower than those expected from ionizing radiation.
+Additionally, we establish thresholds on the maximum number of zeros allowed in the pre-trigger region.
+In this region, the number of zeros should be consistent with those predicted by the qubit decay.
+A significantly larger number of zeros would indicate transient noise that could induce false triggers and has to be discarded from the analysis.
+Therefore, we calculate the Binomial probability of observing a specific number of zeros in the pre-trigger region and apply a cut for probabilities below 1%.
+Detailed thresholds used for each dataset are provided in Supplementary Section F.
+IV Results and Discussion
+The distributions of the number of zeros in the signal region for the triggered events are depicted in Fig.
 4.
-We then computed several quantities to characterize each pulse, including the value of the maximum, the length (defined as the number of points above threshold), the position of the maximum, as well as the value and root-mean-square (RMS) of the baseline before the pulse.
-These parameters were used to discard non-physical pulses (see orange trace in Fig.
-4), such as those occurring on the tail of a preceding pulse or with a maximum too far from the trigger position.
-Consistency in data selection was ensured by applying the same criteria across all runs.
-IV Results and Discussion
-The histogram of Max of the selected pulses is shown in Fig.
+These distributions correspond to background runs conducted at the fully shielded facility at LNGS without any Thorium source and at FNAL, where qubits were entirely exposed to cosmic rays and gamma radiation in the laboratory environment.
+It can be observed that the histograms of Q1 and Q3, which have similar values, also exhibit similar shapes.
+This observation holds true for Q2 and Q4 as well.
+As described in the previous section, to identify radiation-induced events, we reduce qubit spontaneous decay events to negligible rates by setting a threshold on the minimum number of zeros in the signal region.
+For example, for the LNGS–Q1 and FNAL–Q3 datasets, which feature a similar , we required at least 23 zeros (non-consecutive) in order to have a negligible rate from qubit spontaneous decay.
+In contrast, for FNAL–Q2 and FNAL–Q4, we required at least 18 zeros due to the lower .
+We emphasize that such a stringent requirement discards the vast majority of the events (see Fig.
+4).
+Additionally, a maximum number of zeros in the pre-trigger region is set to mitigate noisy periods.
+This number ranges from a minimum of 10, for datasets with lower to a maximum of 15 for datasets with higher .
+The cuts on the number of zeros in the pre-trigger and in the signal regions for all the datasets are reported in Table S4 of the Supplement.
+Table 4 summarizes the results obtained from the background runs conducted at LNGS and FNAL.
+The three qubits operated on the same chip at FNAL show rates ranging from 3 to 2510-3 events/sec, in contrast to what we would expect if the selected events were dominated by radioactivity.
+In this case, indeed, the rates should have been similar.
+Moreover, the rate measured in the fully shielded facility at LNGS is comparable with the one measured at FNAL.
+If the obtained values were ascribed to radioactivity, the LNGS rate should have been an order of magnitude smaller than the FNAL one.
+These results suggest that modern transmons are primarily affected by sources of noise other than radiation, contributing significantly to the reduction in T1 at short timescales.
+This fact has also been corroborated by another recent study [38], where cosmic rays were not identified as a major contributor to the most frequent correlated errors (among four qubits).
+Finally, we observe that the measured rates are lower than those expected from simulations of ambient radiation.
+This discrepancy suggests that the event selection substantially reduces the efficiency of radiation-induced events.
+Datasets | Measured Rate | Radioactivity Contribution |
+[events/sec] | [events/sec] | |
+Shielded chip | (2.9 1.3) | (4.0 0.6) |
+FNAL Q2 | (3.2 0.7) | |
+FNAL Q3 | (4.7 0.9) | (57 3) |
+FNAL Q4 | (25 4) |
+To better investigate the detection efficiency for radiation-induced events, at LNGS, we exposed the chip to Thorium radioactive sources with increasing activity levels.
+Several commercially available Thorium rods with known activity were inserted inside the copper shield and outside the outer vacuum chamber of the fridge for this study.
+Data were analyzed using the same protocol used for background runs and reported in Fig.
 5.
-The single qubit measured in the fully shielded facility at LNGS features a rate of events/sec (solid gray trace).
-This is obtained by using the measurement time of 28 min, which is selected to filter periods with some unwanted excitation of the state.
-This shorter time causes larger statistical uncertainty.
-We repeated the analysis including longer intervals (in which state was sometimes excited) in order to increase the statistics and obtained events/sec (hatched gray trace).
-This factor two discrepancy can be considered as the systematic uncertainty on the result.
-The observed background event rate is about 2.5 to 5 times higher than the predicted rate from simulations, which is events/sec.
-This discrepancy suggests the presence of other sources of decoherence that resemble quasiparticle burst events but cannot be ascribed to radioactive sources.
-Figure 5 clearly demonstrates that exposing the chip to a radioactive source results in an excess of events.
-The rate of these events increases linearly with the intensity of the source itself.
-Next, we compare the triggered and simulated rates for the various Thorium sources in Fig.
+It demonstrates that exposing the chip to a controlled radioactive source results in an excess of events with high number of zeros in the signal region compared to the “shielded” configuration.
+The rate of these events increases linearly with the intensity of the source itself as visible in Fig.
 6.
+Next, we compare the triggered and simulated rates for the various Thorium sources.
 The slope parameter p1 of the linear fit in Fig.
-6 shows that setting the threshold at 1 probability allows us to reach a detection efficiency of .
-This result indicates that qubits are sensitive to laboratory gamma radioactivity, despite the fact that gamma rays deposit a lower amount of energy compared to cosmic rays [22, 33].
-This is not a significant concern for the development of quantum processors, as minimal passive shielding around the facility can largely mitigate this radioactive source [39, 50, 51].
-Conversely, the absence of a high energy threshold for gamma detection presents interesting opportunities for using qubits as particle detectors [52].
-While the primary objective of this work was not to operate a qubit as a particle detector, it is evident that optimizing the threshold or trigger algorithm, along with employing more refined pulse reconstruction algorithms (such as matched filter algorithms), could significantly improve detection efficiency.
-However, from the linear fit the intercept p0 results in () events/second.
-This parameter can be interpreted as the expected rate in the absence of radioactive interactions (if the qubit response to the source is fully linear) and exceeds the one measured in the “background” run.
-Future work will investigate this discrepancy further.
-We deepened the study of the effect of radioactivity by investigating a possible activation of two-level-systems (TLS) when deploying the gamma source.
-As pointed out in Ref.
-23, the ionizing effect of radiation on the chip and the consecutive charge recombination could locally affect the distribution of charges and the electric field that they produce, to which TLSs are sensitive.
-As a result, the TLSs’ frequencies could shift, getting into resonance or out of resonance with the qubit.
-This would cause a fluctuation of the qubit T1 lasting longer than the recombination time of the quasiparticles produced by an impact.
-We analyzed the data obtained from the measurements done at LNGS when exposing the qubit to the 161 kBq Thorium source.
-We selected pulses with very high maxima (), so to discard possible noise events.
-For each pulse we calculated the average values of approximately 10 ms before the trigger and tens of milliseconds after it, when the quasiparticle poisoning was already over.
-A change in the qubit T1 due to TLS dynamics would result in an average value of after the pulse different from what obtained before the trigger.
-We made the histogram of for radioactivity impacts and compared it with what obtained by selecting noise events instead (Fig.
-7).
-We didn’t observe significant differences between the two histograms.
-To double check this result we exposed our qubit to a Cesium source which activity was an order of magnitude higher than the Th source one.
-Also in that case, no significant differences between radioactivity events and noise events were observed.
-This analysis showed that the fluctuations in the qubit T1 correlated to radioactivity impacts were of the same magnitude of what caused by other sources of noise.
-Given also the high rate of noise events that we observed, we concluded that ionizing radiation does not play a major role in undermining the T1 stability of the qubit we studied.
-Finally, we investigated whether radioactive impacts could excite the qubit from to state.
-To this aim, we performed the same study described above, but with a reversed measurement protocol: initializing the qubit in the state instead of , and monitoring the qubit exposed to a strong gamma source.
-Interestingly, exposing the qubit to gamma radiation did not result in an excess of events as demonstrated in Fig.
-10.
-This observation is consistent with what was previously reported in Ref.
-20, confirming that quasiparticles generated due to radiation do not deliver energy to transmon qubits.
-Finally, we conducted a similar fast decay protocol study using multiple “twin-chips” operated at the above ground facilities at SQMS Fermilab.
-Here qubits are fully unshielded from cosmic and gamma radiation.
-The results showed consistency with those obtained at LNGS, as we obtained qualitatively similar distributions in all the pulse parameters defined for our study.
-The distribution of Max in particular, is reported in Figure 8.
-The histograms do not reveal any distinct behavior of the chip in the measurements conducted deep underground at LNGS compared to those above-ground at FNAL.
-The protocol applied to three qubits at FNAL under ambient radiation conditions showed varied responses: similar (Q4), reduced (Q2), and increased (Q3) rates when compared to the response of the maximally shielded qubit at LNGS.
-These variations observed in the rates for different qubits measured at FNAL compared to the LNGS background one indicate that modern transmons are dominated by other sources of noise and radiation does not play a major role in the drastic drop in T1 at short time scales.
-This fact has also been corroborated by another recent study [41] where cosmic rays are not found to be a strong contributor to the most frequent correlated error (four qubits).
-Even though we did not experimentally observe a measurable effect of environmental radiation induced errors, we can utilize the simulated rates to give a cap estimate for the relevant timescales at which radiation would start to affect quantum computation.
+6 shows that the aggressive data selection leads to a detection efficiency of , where the uncertainty is statistic only.
+The systematic uncertainty is much larger, as the signal efficiency depends on the cuts used for data selection, that vary across the datasets because of the differing and noise conditions.
+Therefore, this result should be considered qualitative.
+Correcting the predicted rate at LNGS (410-3 event/sec) by this efficiency results in an expected rate from radioactivity of 0.3210-3 event/sec, an order of magnitude smaller than the measured one.
+This further supports the hypothesis that the measured rates are not dominated by radiation-induced events.
+Despite the low detection efficiency, our findings indicate that qubits are sensitive to laboratory gamma radioactivity, even though gamma rays deposit less energy compared to cosmic rays [22, 34].
+While the primary objective of this work was not to operate a qubit as a particle detector, it is evident that optimizing the thresholds or trigger algorithms, along with employing more refined pulse reconstruction techniques (such as matched filter algorithms), could significantly improve detection efficiency.
+Although we did not experimentally observe a measurable effect of environmental radiation-induced errors, we can utilize the simulated rates to provide an upper estimate for the relevant timescales at which radiation would start to affect quantum computation through information loss occurring on individual qubits.
 The probability to have at least one event in a time window T is given by , where is the rate of impacts.
@@
 Note that the computation may include multiple runs of an algorithm and not necessarily require to be a single run lasting for T.
-At a fully shielded environment like LNGS, due to much lower expected rate, we have instead (1%) if (2.5 s).
-The available computation period increases if higher error probability is tolerated.
+At a fully shielded environment like LNGS, due to a much lower expected rate, we have instead (1%) if (2.5 s).
+The available computation period increases if a higher error probability is tolerated.
 Figure 7 shows available computational time as a function of radiation impact rates around values estimated for the two cases.
@@
 In other words, when a calculation is repeated times, one result will be corrupted due to radiation.
-Most of the modern algorithms running on quantum hardware like variational quantum eigensolver or quantum approximate optimization take less than a milli-second for a single iteration and require a large number of repetitions ().
-Therefore, we can safely conclude that radiation has negligible effect on modern transmon-based quantum computing platforms even if error-correction fails due to an impact.
+Most of the modern algorithms running on quantum hardware like variational quantum eigensolver or quantum approximate optimization take less than a millisecond for a single iteration and require a large number of repetitions ().
+Therefore, we can safely conclude that radiation has a negligible effect on modern transmon-based quantum computing platforms even if error correction fails due to an impact.
 V Conclusions
-We have performed a comparative study of how radiation may affect transmon qubits hosted at two laboratories with very distinct radiation environments.
-The regular T1 experiments performed on the same qubit showed no difference in qubit lifetimes of the order of 100 microseconds.
-The experiments with controlled Thorium sources conducted at the deep underground Gran Sasso laboratory demonstrate the possibility of achieving a good detection efficiency, opening the possibility of using transmons as a particle detector.
-However, the background rate observed with no active sources at LNGS show higher event rates compared to the simulated values, indicating that other noise sources produce signatures similar to those expected or conjectured to be coming from radiation.
-The same protocol when applied to multiple similar qubits in Fermilab (above-ground and with no lead shield) with ambient radiation, shows both increased and reduced rates compared to the response of the maximally shielded qubit at LNGS.
-These results suggest that these radiation-like events lasting milliseconds are generated by different sources, and that radiation is not a major source of errors in modern transmon qubits.
-Based on the experimental and simulated results we find that quantum computation will be affected at a maximum of a percent level due to radiation inflicted single-qubit errors when computational time reaches more than 170 ms at fully unshielded facilities and 2.5 s at LNGS.
-We found no correlation of radiation events to TLS scrambling or qubit excitation.
-Future experiments will aim at in-depth understanding of the various error sources which are currently dominating over the radiation induced effects.
-This will involve further investigations based on fast decay protocols testing different qubit geometries, materials and synchronized measurements across same and different chips.
-SQMS is focused on enhancing the T1 of transmon qubits and so future measurements will investigate the behavior of qubits with a wide distribution of coherence performance ranging from tens of microseconds to approaching milliseconds.
-Advancing this knowledge will be critical for the use of superconducting qubits for quantum computation and sensing.
+We have conducted a comparative study to assess the impact of radiation on transmon qubits at two laboratories with very different radiation environments.
+Regular T1 experiments performed on the same qubit show no significant difference in qubit lifetimes of the order of 80 microseconds.
+In order to study the time dynamics at short timescales, we employed a fast decay detection protocol utilizing active reset.
+Multiple similar qubits measured with this technique at Fermilab (above ground and without lead or copper shieldings) showed rates comparable to those from the maximally shielded qubit at the deep-underground Gran Sasso Laboratory.
+These findings suggest that the radiation-like events observed, which last for milliseconds, originate from different sources, and indicate that radiation is not a major source of errors in modern transmon qubits.
+Experiments with controlled Thorium sources at LNGS demonstrate the potential for detecting particle impacts, opening the possibility of using transmons as particle detectors.
+However, the observed rates are lower than those predicted by simulations, indicating the need for more refined trigger algorithms and/or analysis strategies.
+Based on both experimental and simulated results, we estimate that radiation-induced single-qubit errors will affect quantum computation by at most a few percent, with computational times up to 170 ms at fully unshielded facilities and 2.5 s at LNGS.
+Future experiments will focus on a more detailed understanding of the various error sources currently overshadowing radiation-induced effects.
+This will involve further investigations using fast decay detection protocols, exploring different qubit geometries, materials, and synchronized measurements across the same and different chips.
+The SQMS Center aims to enhance the T1 of transmon qubits, and future studies will examine the performance of qubits with a broad range of coherence times, from tens of microseconds to approaching milliseconds.
+Advancing this knowledge is crucial for the effective use of superconducting qubits in quantum computation and sensing.
 Acknowledgements.
@@
 We are also grateful to the LNGS Computing and Network Service for computing resources and support on U-LITE cluster at LNGS.
+We thank S.
+Nisi for the ICP-MS measurements of the PCB, C.
+Tomei and G.
+D’Imperio for the maintenance of the simulation software, G.
+Catelani for the useful discussion.
 We are grateful to A.
@@
 III and Grzegorz Tatkowski for providing support on experimental preparations and cryogenic operations at FNAL.
-T.R.
-thanks Xinyuan You for useful discussions.
 The work was supported by the U.S.
 Department of Energy, Office of Science, National Quantum Information Science Research Centers, Superconducting Quantum Materials and Systems (SQMS) Center under the contract No.
-DE-AC02-07CH11359, by the Italian Ministry of Foreign Affairs and International Cooperation, grant number US23GR09, by the Italian Ministry of Research under the PRIN Contract No.
+DE-AC02-07CH11359, by the Italian Ministry of Foreign Affairs and International Cooperation, grant number US23GR09, by the PNRR MUR project CN00000013-ICSC, by the Italian Ministry of Research under the PRIN Contract No.
 2020h5l338 (“Thin films and radioactivity mitigation to enhance superconducting quantum processors and low temperature particle detectors”) and by the Italian Ministry of Research under the PRIN Contract No.
@@
 - McRae et al. [2020] C. R. H. McRae et al., Materials loss measurements using superconducting microwave resonators, Review of Scientific Instruments 91, 091101 (2020).
+- Bratrud et al.
+[2024] G.
+Bratrud, S.
+Lewis, K.
+Anyang, A.
+Cesaní, T.
+Dyson, H.
+Magoon, D.
+Sabhari, G.
+Spahn, G.
+Wagner, R.
+Gualtieri, et al., First measurement of correlated charge noise in superconducting qubits at an underground facility, arXiv preprint arXiv:2405.04642 (2024).
 - Gusenkova et al.
@@
 Plourde, Phonon downconversion to suppress correlated errors in superconducting qubits, Nature Communications 13, 6425 (2022).
+- Harrington et al.
+[2024] P.
+M.
+Harrington, M.
+Li, M.
+Hays, W.
+V.
+D.
+Pontseele, D.
+Mayer, H.
+D.
+Pinckney, F.
+Contipelli, M.
+Gingras, B.
+M.
+Niedzielski, H.
+Stickler, J.
+L.
+Yoder, M.
+E.
+Schwartz, J.
+A.
+Grover, K.
+Serniak, W.
+D.
+Oliver, and J.
+A.
+Formaggio, Synchronous detection of cosmic rays and correlated errors in superconducting qubit arrays (2024), arXiv:2402.03208 [quant-ph] .
 - Le et al. [2021] D. T. Le et al., Operating a passive on-chip superconducting circulator: Device control and quasiparticle effects, Phys. Rev. Research 3, 043211 (2021).
@@
 - Orrell and Loer [2021] J. L. Orrell and B. Loer, Sensor-assisted fault mitigation in quantum computation, Phys. Rev. Applied 16, 024025 (2021).
-- Harrington et al.
-[2024] P.
-M.
-Harrington, M.
-Li, M.
-Hays, W.
+- Mariani et al. [2023] A. Mariani et al., Mitigation of cosmic rays-induced errors in superconducting quantum processors, IEEEQCE23 proceedings (2023).
+- Yelton et al.
+[2024] E.
+Yelton, C.
+P.
+Larson, V.
+Iaia, K.
+Dodge, G.
+L.
+Magna, P.
+G.
+Baity, I.
 V.
-D.
-Pontseele, D.
-Mayer, H.
-D.
-Pinckney, F.
-Contipelli, M.
-Gingras, B.
-M.
-Niedzielski, H.
-Stickler, J.
+Pechenezhskiy, R.
+McDermott, N.
+Kurinsky, G.
+Catelani, and B.
 L.
-Yoder, M.
-E.
-Schwartz, J.
-A.
-Grover, K.
-Serniak, W.
-D.
-Oliver, and J.
-A.
-Formaggio, Synchronous detection of cosmic rays and correlated errors in superconducting qubit arrays (2024), arXiv:2402.03208 [quant-ph] .
-- Mariani et al. [2023] A. Mariani et al., Mitigation of cosmic rays-induced errors in superconducting quantum processors, IEEEQCE23 proceedings (2023).
+T.
+Plourde, Modeling phonon-mediated quasiparticle poisoning in superconducting qubit arrays (2024), arXiv:2402.15471 [quant-ph] .
+- Martinez et al.
+[2019] M.
+Martinez, L.
+Cardani, N.
+Casali, A.
+Cruciani, G.
+Pettinari, and M.
+Vignati, Measurements and simulations of athermal phonon transmission from silicon absorbers to aluminum sensors, Phys.
+Rev.
+Appl.
+11, 064025 (2019).
 - Bal et al.
@@
 Grassellino, Systematic improvements in transmon qubit coherence enabled by niobium surface encapsulation, npj Quantum Information 10, 43 (2024).
-- Yelton et al.
-[2024] E.
-Yelton, C.
-P.
-Larson, V.
-Iaia, K.
-Dodge, G.
-L.
-Magna, P.
-G.
-Baity, I.
-V.
-Pechenezhskiy, R.
-McDermott, N.
-Kurinsky, G.
-Catelani, and B.
-L.
-T.
-Plourde, Modeling phonon-mediated quasiparticle poisoning in superconducting qubit arrays (2024), arXiv:2402.15471 [quant-ph] .
-- Martinez et al.
-[2019] M.
-Martinez, L.
-Cardani, N.
-Casali, A.
-Cruciani, G.
-Pettinari, and M.
-Vignati, Measurements and simulations of athermal phonon transmission from silicon absorbers to aluminum sensors, Phys.
-Rev.
-Appl.
-11, 064025 (2019).
 - Agostinelli et al. [2003] S. Agostinelli et al., GEANT4—a simulation toolkit, Nucl. Instrum. Methods. Phys. Res. A 506, 250 (2003).
@@
 Schuster, The qick (quantum instrumentation control kit): Readout and control for qubits and detectors (2022), arXiv:2110.00557 [quant-ph] .
-- Fowler et al.
-[2024] J.
-W.
-Fowler, P.
-Szypryt, R.
-Bunker, E.
-R.
-Edwards, I.
-F.
-Florang, J.
-Gao, A.
-Giachero, S.
-F.
-Hoogerheide, B.
-Loer, H.
-P.
-Mumm, N.
-Nakamura, G.
-C.
-O’Neil, J.
-L.
-Orrell, E.
-M.
-Scott, J.
-Stevens, D.
-S.
-Swetz, B.
-A.
-VanDevender, M.
-Vissers, and J.
-N.
-Ullom, Spectroscopic measurements and models of energy deposition in the substrate of quantum circuits by natural ionizing radiation (2024), arXiv:2404.10866 .
-- Loer et al.
-[2024] B.
-Loer, P.
-M.
-Harrington, B.
-Archambault, E.
-Fuller, B.
-Pierson, I.
-Arnquist, K.
-Harouaka, T.
-D.
-Schlieder, D.
-K.
-Kim, A.
+- Adari et al. [2022] P. Adari et al., Excess workshop: Descriptions of rising low-energy spectra, SciPost Physics Proceedings 10.21468/scipostphysproc.9.001 (2022).
+- Anthony-Petersen et al. [2022] R. Anthony-Petersen et al., A stress induced source of phonon bursts and quasiparticle poisoning (2022), arXiv:2208.02790 [physics.ins-det] .
+- Angloher et al. [2023] G. Angloher et al., Latest observations on the low energy excess in cresst-iii, SciPost Physics Proceedings 10.21468/scipostphysproc.12.013 (2023).
+- Armengaud et al. [2016] E. Armengaud et al., Constraints on low-mass wimps from the edelweiss-iii dark matter search, Journal of Cosmology and Astroparticle Physics 2016 (05), 019–019.
+- Romani [2024] R. K. Romani, Aluminium relaxation as the source of excess low energy events in low threshold calorimeters (2024), arXiv:2406.15425 [physics.ins-det] .
+- Kono et al.
+[2024] S.
+Kono, J.
+Pan, M.
+Chegnizadeh, X.
+Wang, A.
+Youssefi, M.
+Scigliuzzo, and T.
 J.
-Melville, B.
-M.
-Niedzielski, J.
-K.
-Yoder, K.
-Serniak, W.
-D.
-Oliver, J.
-L.
-Orrell, R.
-Bunker, B.
-A.
-VanDevender, and M.
-Warner, Abatement of ionizing radiation for superconducting quantum devices (2024), arXiv:2403.01032 .
-- Linehan et al.
-[2024] R.
-Linehan, I.
-Hernandez, D.
-J.
-Temples, S.
-Q.
-Dang, D.
-Baxter, L.
-Hsu, E.
-Figueroa-Feliciano, R.
-Khatiwada, K.
-Anyang, D.
-Bowring, G.
-Bratrud, G.
-Cancelo, A.
-Chou, R.
-Gualtieri, K.
-Stifter, and S.
-Sussman, Estimating the energy threshold of phonon-mediated superconducting qubit detectors operated in an energy-relaxation sensing scheme (2024), arXiv:2404.04423 .
+Kippenberg, Mechanically induced correlated errors on superconducting qubits with relaxation times exceeding 0.4 ms, Nature Communications 15, 3950 (2024).
+- Verkerke and Kirkby [2003] V. Verkerke and D. Kirkby, The roofit toolkit for data modeling (2003), arXiv:0306116 .
 Author Contributions
@@
 The substrate preparation, base-layer (Nb) deposition and Josephson junction evaporation are similar to the fabrication methods followed in Ref.
-43.
-In order to prevent oxidation of the Niobium, a capping layer of thickness nm is deposited using the same sputtering tool without breaking the vacuum.
+46.
+In order to prevent oxidation of the Niobium, a nm thick capping layer is deposited using the same sputtering tool without breaking the vacuum.
 Gold and Tantalum are used as capping layers for devices Q1 and Q2-Q4 respectively (on two different chips).
@@
 Readout frequency | 7206.8 | 7055.0 | 6886.5 | 6714.5 | MHz |
-Qubit pulse length | 0.150 | 0.091 | 0.124 | 0.160 | s |
-Qubit average T1 | 84 | 141 | 131 | 214 | s |
-Readout pulse length | 4.5 | 3.8 | 4.0 | 8.0 | s |
-Waiting period | 5.0 | 10.0 | 5.0 | 5.0 | s |
-Cooldown period | 50.0 | 70.0 | 70.0 | 10.0 | s |
-One iteration period | 64.550 | 87.929 | 84.324 | 31.660 | s |
+Qubit pulse length | 0.195 | 0.091 | 0.124 | 0.160 | s |
+Readout pulse length | 6.0 | 3.8 | 4.0 | 8.0 | s |
 V.2 Measurement Setup
@@
 The input line is attenuated by 60-70 dB distributed between the 4-K and mixing chamber stages.
-Copper-body 20 dB attenuators are used at the base plate plate for better thermalization.
+Copper-body 20 dB attenuators are used at the base plate for better thermalization.
 At LNGS, the input line is also filtered through a 8 GHz reflective low-pass filter and a strong infrared filter (QMC-CRYOIRF-002) to reduce high-frequency noise.
@@
 f.
-A two layers Permimphy magnetic shielding (Outer layer: 55.6 cm (OD) 91.3 cm, 1.5 mm-thick; Inner layer: 51 cm (ID) 89 cm, 1.5 mm-thick);
+A two layers Permimphy magnetic shielding (Outer layer: 55.6 cm (OD) 91.3 cm (height), 1.5 mm-thick; Inner layer: 51 cm (ID) 89 cm (height), 1.5 mm-thick);
 g.
-A 48.1 cm (OD) 119.5 cm simplified version of the cryostat with internal vessels;
+A 48.1 cm (OD) 119.5 cm (height) simplified version of the cryostat with internal vessels;
 h.
@@
 The event rates for each radiation source are reported in Table 1.
-Notably, the overall rate is expected to be approximately an order of magnitude lower at INFN-LNGS (4 @INFN-lNGS vs.
-57 @FNAL), where the dominant contribution is from the PCBs.
+Notably, the overall rate is expected to be approximately an order of magnitude lower at INFN-LNGS (4 event/sec vs.
+57 event/sec @FNAL), where the dominant contribution is from the PCBs.
 To investigate the rate in the presence of the Thorium radioactive sources used at INFN-LNGS, we generate radioactive decays from a point-like 232Th radioactive source positioned between the external magnetic shield and the copper bars, approximately at the same height of the chip.
 For this simulation, we consider only gamma emitters in the chain (228Ra, 228Ac, 212Pb, 212Bi and 208Tl), since and particles cannot reach the chip as they are shielded by the cryostat.
-As in the previous case, we determine the rate of interactions in the chip by scaling the number of recorded events to the activities reported in Table S3, which correspond to the Thorium source activities employed in the measurements.
+As in the previous case, we determine the rate of interactions in the chip by scaling the number of recorded events to the activities reported in Table S3, which correspond to the Thorium source activities employed in the measurements (and previously characterized using High Purity Germanium spectroscopy).
 The rate of events in the chip, plotted as a function of the deposited energy for the various activities, is illustrated in Fig.
@@
 V.4 Qubit State Discrimination
-The I/Q values (in-phase and quadrature-phase signals) obtained for each measurement depend on the state of the qubit.
-When a large counts of measurements are plotted on the I-Q plane, we usually observe two blobs corresponding to and .
-However, we sometimes observe signals corresponding the transmon being in the second excited state as shown in Fig.
-S6 and we discard those measurements in the analysis.
-The blobs corresponding to and are rotated in order to have the centers at the same value of Q and projected onto the I axis.
-In this way it is possible to discriminate between the two states by putting a threshold value only on one (I) quadrature.
-The populations of the states have been estimated by doing a simultaneous fit on the two quadratures with the RooFit toolkit.
-The fit function used was a sum of 2D gaussians, one for each blob we were observing.
-The data was divided into traces of approx.
-1 minute duration, and for each of them the simultaneous fit was performed.
-In this way we were able to identify periods in which the qubit was showing a high population of the state and reject them.
-V.5 Data Analysis
-In our analysis we selected events where the probability of observing a certain number of zeros and ones, calculated according to the Binomial distribution, was below a predefined threshold.
-Specifically, if the probability of observing such a series of zeros and ones fell below 1%, corresponding to an absolute logarithm () greater than 4.6, the software trigger was fired, indicating the presence of a potential signal.
-Once triggered, these data wedre recorded for further analysis.
-By selecting pulses with , we observe a trigger rate of 5–6 events/second in the runs performed with the radioactive sources, significantly higher than the expected rate from radioactivity (0.1–0.4 events per second depending on the source activity), suggesting that the majority of triggers are due to noise.
-To address this, we computed additional variables to refine our data selection.
-Specifically, we calculated:
+The I/Q values (in-phase and quadrature-phase signals) obtained from each measurement reflect the state of the qubit.
+When a large number of measurements are plotted on the I-Q plane, two clusters, corresponding to the states (ground) and (excited), are observed.
+In our data we observe also signals corresponding to the qubit being in the second excited state and, occasionally, in the third excited state .
+To ensure accuracy we discard measurements where the populations of and are too high.
+To identify and exclude periods where the qubit exhibits high populations of the and states, we fit the clouds with 2D Gaussian functions, one for each cluster observed.
+In our experiment the data are divided into traces of measurements.
+The fit is performed on each trace.
+To fit the clouds we use the RooFit toolkit [57].
+The fit is performed by projecting the clouds along the two quadratures and simultaneously fitting both projections.
+The fit function for each projection is given by , where is a Gaussian centered in and with standard deviation .
+Each Gaussian is normalized such that corresponds to the number of measurements associated with that state.
+The population of a state is given then by .
+Fig.
+S6 shows the clouds produced by a single trace and the result of the simultaneous fit used to estimate state populations.
+The clouds corresponding to and are orientated such that their centers are aligned along the same Q value.
+In this way it is possible to discriminate between the two states by setting a threshold only on the I axis, simplifying state identification.
+V.5 “Effective” calculation
+To calculate the “effective” relaxation time over a trace of points, we perform the following steps:
 1.
-The value and RMS of the baseline before the pulse;
+We count the number of times the qubit is in the excited state () OVER the total number of measurements ;
 2.
-The maximum value of ;
-3.
-The pulse length, defined as the number of points above threshold;
-4.
-The position of the maximum.
-The baseline and its RMS were computed using the first 100 points preceding the trigger.
-We then removed pulses occurring on the tail of a preceding pulse or those with a “noisy” baseline (see Fig.
-4) by requiring a baseline value lower or equal to its mean value + 3.
-Additionally, we ensured that the position of the maximum was close to the trigger to avoid re-triggering of the same pulse.
-Since radioactivity typically results in long pulses (at least 1 ms), we further refined our selection by choosing pulses with length greater than 50 points, corresponding to approximately 3.5 ms under our data-taking conditions.
-Finally, we conducted a study to assess the suitability of the window size used for computing the Binomial probability.
-Our initial choice of 50 points stemmed from the analysis of typical events detected in calibration runs.
-By averaging tens of these events, we determined that the qubit returned to equilibrium with a characteristic time constant () of 0.78 ms.
-To ensure complete recovery of the qubit, we adopted a window size corresponding to 5, or approximately 4 ms, which translates to 56 points given our sampling frequency of about 0.07 ms-1.
-We subsequently performed the analysis using window sizes of 20, 30, 70, and 90 points.
-Our findings confirmed that, for all these window sizes, the detection efficiency could not reach values above 70.
-Indeed, when the window size is too small, only a fraction of the signal can be captured within the window.
-Conversely, if the window size exceeds the signal length, the number of zeros associated with random qubit de-excitation becomes significant, thus deteriorating the signal-to-noise ratio.
+We calculate as .
+For each dataset, this process is repeated for each segment of points to monitor changes over time, and then we calculate the average , which is used to estimate the probability of the qubit being in state .
+This method ensures accurate determination of the qubit’s relaxation time by averaging over a large number of measurements.
+V.6 Threshold values
+To distinguish events caused by ionizing radiation from those due to qubit spontaneous decay, we employed a consistent analysis approach across all datasets collected at both LNGS and FNAL.
+Our method involves calculating the probability of observing more than a certain number of zeros within a 60-point signal window using the CDF of the Binomial distribution.
+This allows us to establish dataset-specific thresholds aimed at minimizing the contribution from qubit spontaneous decay while potentially maximizing sensitivity to radiation-induced events.
+Specifically, these thresholds are set to target spontaneous qubit decay rates approximately an order of magnitude lower than those expected from ambient radiation sources.
+Additionally, to mitigate noise-induced triggers, a separate threshold is applied to the pre-trigger region, ensuring robustness in our event selection criteria.
+Table S4 summarizes the specific thresholds applied to each dataset, alongside the expected rates from qubit spontaneous decay (calculated using these thresholds), the expected rates from ionizing radiation, and the measured rates.
+Datasets | Noise | Signal | Qubit Spontaneous | Ionizing Radiation | Measured |
+Threshold | Threshold | Decay Rate | Rate | Rate | |
+[# of zeros] | [# of zeros] | [events/sec] | [events/sec] | [events/sec] | |
+Shielded chip | 15 | 23 | 0.0001 | 0.0040 0.0006 | 0.003 0.001 |
+44 kBq Th Source | 14 | 24 | 0.001 | 0.12 0.01 | 0.032 0.002 |
+76 kBq Th Source | 14 | 23 | 0.002 | 0.20 0.02 | 0.036 0.004 |
+125 kBq Th Source | 15 | 24 | 0.004 | 0.34 0.03 | 0.052 0.006 |
+161 kBq Th Source | 15 | 24 | 0.002 | 0.43 0.04 | 0.056 0.004 |
+FNAL Q2 | 10 | 18 | 0.002 | 0.057 0.003 | 0.003 0.001 |
+FNAL Q3 | 14 | 23 | 0.002 | 0.057 0.003 | 0.005 0.001 |
+FNAL Q4 | 10 | 18 | 0.001 | 0.057 0.003 | 0.025 0.004 |